### v2

# Large language model psychometrics: A systematic review of evaluation, validation, and enhancement

Abstract
The advancement of large language models (LLMs) has outpaced traditional evaluation methodologies. This progress presents novel challenges, such as measuring human-like psychological constructs, moving beyond static and task-specific benchmarks, and establishing human-centered evaluation. These challenges intersect with psychometrics, the science of quantifying the intangible aspects of human psychology, such as personality, values, and intelligence. This review paper introduces and synthesizes the emerging interdisciplinary field of LLM Psychometrics, which leverages psychometric instruments, theories, and principles to evaluate, understand, and enhance LLMs. The reviewed literature systematically shapes benchmarking principles, broadens evaluation scopes, refines methodologies, validates results, and advances LLM capabilities. Diverse perspectives are integrated to provide a structured framework for researchers across disciplines, enabling a more comprehensive understanding of this nascent field. Ultimately, the review provides actionable insights for developing future evaluation paradigms that align with human-level AI and promote the advancement of human-centered AI systems for societal benefit. A curated repository of LLM psychometric resources is available at https://github.com/valuebyte-ai/Awesome-LLM-Psychometrics.
1 Introduction
The advent of large language models (LLMs) represents a transformative breakthrough in AI. These systems exhibit general-purpose capabilities spanning diverse domains [Bubeck et al., 2023], with particular proficiency in natural language understanding and generation [Demszky et al., 2023, Grossmann et al., 2023, Ziems et al., 2024, Gu et al., 2024]. They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [OpenAI, 2025] and search engines [Wang et al., 2024c] to high-stakes domains such as healthcare [Singhal et al., 2023], education [Milano et al., 2023], and scientific discovery [Romera-Paredes et al., 2024, Ye et al., 2024]. Their increasing dominance has exposed a fundamental, pressing scientific challenge: how can we rigorously evaluate these AI systems that transcend traditional benchmarks of biological or algorithmic intelligence?
Traditional AI evaluation has relied on curating task-specific datasets, annotating ground-truth labels with human input, running models on these datasets, and assessing performance using predefined metrics. However, LLMs have triggered an evaluation crisis, as their versatile capabilities and human-like behaviors exceed what traditional benchmarks can measure. Novel challenges include but are not limited to (1) evaluating psychological constructs like personality, values, and cognitive biases; (2) obsolescence of static benchmarks due to rapid LLM development and training data contamination; (3) compromised robustness and validity due to LLMs’ prompt- and context-sensitivity; (4) requiring human-centered evaluation approaches; and (5) expanding evaluation methodologies in scope and complexity as LLMs integrate into agentic and multimodal systems.
These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology [Pasquali, 2009]. Psychometrics emerged from this timeless pursuit as the scientific study of psychological measurement. It bridges the abstract and the empirical by transforming human traits into quantifiable data, enabling better understanding, prediction, and decision-making in education, business, healthcare, governance, and beyond [Rust and Golombok, 2014]. The intersection of LLMs and psychometrics forms a methodological crucible, forging novel paradigms to better decode and improve machine minds.
This intersection sets the stage for a new research frontier. We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles. This field seeks to quantify, interpret, manipulate, and improve the complex, human-like attributes and behaviors exhibited by LLMs, encompassing both personality constructs (e.g., personality, values, morality, and attitudes) and cognitive constructs (e.g., heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities). Research in LLM Psychometrics applies, extends, and innovates upon scientific methods of psychological measurement for LLMs. Based on the psychometric principles, related research systematically examines the measurement results to ensure scientific rigor. By measuring and elucidating psychological constructs in LLMs, LLM Psychometrics further informs strategies for their targeted enhancement.
Recent research in LLM Psychometrics pioneers in addressing the LLM evaluation crisis. Some studies introduce dynamic and construct-oriented evaluation frameworks that move beyond static, task-specific benchmarks [Zhu et al., 2024a, Hagendorff, 2023]. In parallel, novel methodologies are developed to measure non-cognitive and emergent constructs [Ren et al., 2024, Huang et al., 2023d, Pellert et al., 2024]. Self-adaptive evaluation techniques now allow for the extrapolation of item difficulty and tailoring assessments to model performance [Jiang et al., 2024a, Lalor et al., 2024, Polo et al., 2024]. Drawing from the methodological framework of psychometric validation, research improves the reliability and validity of evaluation protocols [Ye et al., 2025a]. Human-centered evaluation drives aligning model behavior with human values [Wang et al., 2024f, Yao et al., 2025a]. In addition, the scope of evaluation expands to agentic and multimodal systems, further broadening the methodological landscape [Li et al., 2024b, Huang et al., 2024c].
The field of LLM Psychometrics has seen significant growth, evidenced by the proliferation of related research papers. These studies, however, address a variety of psychological constructs, employ diverse methodologies, and utilize distinct validation techniques. The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic communities. Despite this diversity, there is a lack of cohesion among researchers, leading to a fragmentation of insights, particularly between studies focusing on disparate constructs. Consequently, there is a pressing need for a systematic review to synthesize these efforts and facilitate a more integrated understanding of the field.
Figure 1: Overview of this review.
Taxonomy and paper structure.
To bridge the gap, we systematically review LLM Psychometrics across evaluation, validation, and enhancement. The evaluation framework encompasses three core dimensions: the target construct (what to measure), the measurement method (how to measure), and the validation of results (how well do we measure). The psychometric insights not only inform evaluation but also guide the development and refinement of LLMs (how to improve). Accordingly, this review is structured as follows (Fig. 1). § 2 provides an overview of the preliminaries and methodological foundations to facilitate subsequent discussions. § 3 thoroughly contrasts psychometrics and traditional AI benchmarking and reviews how psychometric principles can underpin and reinvent LLM benchmarks. The core measurement framework is detailed in § 4, § 5, and § 6. § 4 delves into the psychological constructs evaluated in LLMs, elucidating the theories employed and summarizing key evaluation findings. In § 5, we scrutinize the psychometric evaluation methodologies applied to LLMs, followed by § 6, which examines the psychometric validation of evaluation results. Beyond evaluation, § 7 introduces strategies for enhancing LLMs through psychometric insights. § 8 discusses emerging trends, challenges, and future directions. Finally, § 9 highlights key ethical considerations, while § 10 concludes the paper.
Scope and related work.
We consider studies to be out of scope if they do not employ psychometric approaches, adhere to psychometric principles, or if they focus solely on scalar performance metrics rather than characterizing behaviors. Readers interested in conventional LLM benchmarking are encouraged to consult surveys on LLM evaluation [Guo et al., 2023b, Chang et al., 2024]. Several related reviews focus on the evaluation of specific constructs in LLMs, such as personality [Wen et al., 2024b, Dong et al., 2025], attitudes and values [Ma et al., 2024a], cultural awareness [Pawar et al., 2024, Adilazuarda et al., 2024], and theory of mind [Sarıtaş et al., 2025, Dong et al., 2025]. Hagendorff [2023], Hagendorff et al. [2024] introduce the notion of machine psychology and review emergent LLM abilities; however, they do not provide comprehensive coverage of the related research, nor do they elaborate on LLM personality constructs, psychometric validation, or enhancement. This paper provides the first systematic review of LLM Psychometrics.
2 Preliminary and methodological foundations
We aim for our survey to be self-contained and accessible to a broad, cross-disciplinary audience. To this end, this section presents the preliminary and methodological foundations that underpin the subsequent discussions.
2.1 Large language models
LLMs are large-scale deep neural networks—essentially complex systems of nonlinear regression equations. An LLM can generate text by predicting the next token (word or subword) in a sequential manner (autoregressive generation), given the preceding context. It does so by modelling a conditional probability distribution over the vocabulary; i.e., the likelihood of each token given the context:
$P(x_{t}|x_{<t})=f(x_{<t};\theta),$ (1)
where $x_{t}$ is the token at time step $t$; $x_{<t}$ is the context preceding $x_{t}$, usually including both user prompts and previously generated tokens; $f$ is the model’s parameterized function; and $\theta$ represents the model parameters. Given $f$, the model generates text by either sampling from this distribution or directly selecting the token with the highest probability. In the former case, hyperparameters such as temperature can be adjusted to control the diversity of generated text. In the latter case, the model is said to use greedy decoding, and the generated text is deterministic. In evaluating LLMs, it is crucial to properly account for the stochasticity of the model.
These models are based primarily on the transformer architecture, a neural network design that employs self-attention mechanisms to capture contextual relationships between words, phrases, and broader linguistic patterns. Modern LLMs typically contain billions of parameters, enabling them to efficiently learn from vast amounts of textual data. During evaluation, if the model has already been exposed to the test items during training, this is referred to as data contamination. In such cases, the model is more likely to exhibit artificially inflated performance or simply reproduce memorized patterns, rather than revealing its true underlying capabilities or traits.
The training process of LLMs is typically divided into two phases: pre-training and post-training. Pre-training is the phase where LLMs learn to predict the next token, given its preceding context, on a large corpus of text data. This process is unsupervised, as the model does not require explicit labels or annotations to learn the underlying patterns in the data. The model processes Internet-scale text data from diverse sources like books, articles, and websites. By repeatedly predicting the next word in a sentence, the model learns the statistical properties of language and gains large-scale world knowledge. Models that have only undergone the pre-training phase are usually referred to as base models. Post-training, or fine-tuning, is the process of adapting the base models to better follow user instructions, align with human values, or specialize in particular tasks. This stage typically involves training the model on a smaller, human-annotated dataset or incorporating human feedback on the quality of the model outputs. Models that have undergone both phases are often referred to as fine-tuned models, instruction-tuned models, or aligned models.
We interact with LLMs using prompts, which are input instructions to the model. For psychometric evaluation, these prompts can naively be the reformatted versions of test items originally designed for humans, adapted for LLMs to answer. When designing prompts, one should consider differences between base and fine-tuned models. Most public-facing LLMs are fine-tuned, so evaluation research primarily focuses on these models for greater practical relevance.
A key emergent capability of LLMs is in-context learning, which allows models to adapt to new tasks or patterns by conditioning on examples or instructions provided within the input context $x_{<t}$, without modifying model parameters. This property can influence LLM performance in psychometric evaluation. For instance, prompting models to reason step-by-step (e.g., Chain-of-Thought prompting [Wei et al., 2022]) can enhance performance on reasoning tasks, while instructing them to role-play may modulate their exhibited personalities and values.
2.2 Psychometrics
Psychometrics, also known as psychological testing, involves the use of tests to measure, understand, or predict behavior by quantifying specific actions or characteristics. These tests rely on samples of behavior, meaning they are not perfect measures and often include errors inherent to sampling. Test items are specific stimuli designed to elicit observable reactions that can be scored or evaluated. Typically, tests are composed of multiple questions or problems as their items, producing explicit data subject to scientific analysis.
A psychological test is a set of items that are designed to measure characteristics of human beings that pertain to behavior [Kaplan and Saccuzzo, 2001]. Behavior measured by tests can be overt (observable actions) or covert (internal thoughts or feelings). Tests may assess past, current, or even predict future behavior. Interpretation of test scores depends on their context within a distribution. Scales are used to relate raw scores to defined distributions, aiding interpretation. In addition, psychological tests can measure traits—enduring tendencies like shyness or determination—and states, which reflect temporary conditions of individuals.
Psychological testing measures individual differences in various constructs, which are abstract psychological attributes or dimensions that help explain and predict behavior. Two primary categories of such constructs are personality constructs and cognitive constructs [Kaplan and Saccuzzo, 2001]. Personality tests focus on an individual’s tendencies and dispositions. These tests measure typical behavior, such as preferences or tendencies to react in certain ways. Cognitive tests evaluate speed, accuracy, or both, with higher scores reflecting better performance.
Two fundamental principles underpin psychometrics: reliability and validity [Raykov and Marcoulides, 2011]. Reliability ensures accuracy, dependability, consistency, or repeatability of the test results. Reliable test results are stable across time, contexts, and raters. Validity confirms the meaningfulness and usefulness of the test results. A valid measure captures the intended construct. Validity is multifaceted; for example, predictive validity might correlate test scores with job performance, while construct validity ensures alignment with theoretical models, such as the Big Five personality traits [Goldberg, 2013].
Other principles include standardization, which provides context to raw scores by comparing individual results to a representative sample, or norm group. Additionally, equivalence and fairness are crucial principles that tests must adhere to. Test bias occurs when test items unintentionally advantage or disadvantage subgroups. Modern psychometrics employs advanced statistical models to identify and revise biased items, ensuring assessments measure the intended construct rather than extraneous factors [Rust and Golombok, 2014].
2.3 Psychometric evaluation of AI before the era of LLMs
The idea of applying psychometrics to AI originated in the early decades of AI [Pellert et al., 2024]. Evans [1964] pioneered work in this area by creating a heuristic program that could solve parts of intelligence tests. Subsequent efforts similarly focused on designing AI systems for cognitive tests [Newell, 1973], with the goal of creating systems capable of handling human tasks. This was conceptually aligned with the development of static, task-centric benchmarks in modern AI research [Hendrycks et al., 2020, Chen et al., 2021, Liang et al., 2022, Srivastava et al., 2022, Lee et al., 2024e]. However, criticisms emerged regarding the absence of "hot cognition" in AI, prompting Simon [1963] to propose incorporating emotional aspects into models. By the early 2000s, the concept of "psychometric AI" was explicitly articulated as the pursuit of systems capable of excelling on all established, validated tests of intelligence and mental ability. These included not only conventional IQ tests but also assessments of artistic and literary creativity, mechanical ability, and beyond [Bringsjord and Schimanski, 2003, Pellert et al., 2024]. It was not until the advent of LLMs that the versatility envisioned for "psychometric AI" began to materialize.
3 Psychometrics for benchmarking principles
3.1 Fundamental differences between psychometrics and AI benchmarking
Benchmarking AI systems superficially resembles psychometrics, particularly Classical Test Theory (CTT) [Crocker and Algina, 1986], as both compile test items to evaluate cognitive capabilities and average the resulting scores. However, closer examination reveals that AI benchmarks differ significantly from modern psychometric approaches [Wang et al., 2023a, Federiakin, 2025]. We outline these key differences in Table 1.
Table 1: Comparison between psychometrics and conventional AI benchmark.
Feature Psychometrics AI benchmark
Core goal To measure psychological constructs, to prove that a test measures as intended (validity evidence), and to understand the construct being measured. To test and compare the task performance of different LLMs. Focuses on ranking models and selecting the best one suited for a specific task.
Philosophy of measurement Construct-oriented. Tends towards a causal approach to measurement, where the measured trait is believed to cause the measurement outcomes. Task-oriented. Leans towards representativism, assuming items exhaust or represent all aspects of the underlying ability.
Target construct Personality and ability. Mostly task-specific abilities.
Construct definition Emphasizes clear and detailed definitions of the construct being measured. Agreement on the construct definition is a byproduct of test development. Often defines constructs implicitly through ad hoc task selection. Construct definitions can be vague.
Development process Systematic and rigorous, often following methods like Evidence-Centered Design (ECD). Can be labor-intensive. Compiles a set of relevant questions or tasks, then performs expert annotation or crowdsourcing to label ground truth answers. Less labor-intensive per item.
Number of items Can vary, but not necessarily large. Focus is on item quality and relevance to the construct. Typically consists of an extensive number of questions to cover various aspects of abilities. Reliability increases with test length.
Sample size Typically requires a larger sample size of test takers for robust statistical modeling. Can be applied to evaluate the performance of a single LLM on the benchmark.
Statistical modeling Employs advanced and various statistical models like Item Response Theory and Factor Analysis to analyze data, estimate latent abilities, and assess model fit. Often relies on simple aggregation methods, such as calculating average accuracy across benchmark tasks.
Result analysis Ensures the reliability, validity, predictive power, and explanatory power of the test through result analysis and statistical modeling. Reliability is likely to be high due to the large number of items. However, validity, predictive power, or explanatory power beyond the target task is not a primary concern.
Psychometrics.
Psychometrics centers on understanding the psychological constructs and ensuring that tests accurately measure the intended constructs. Grounded in a causal measurement philosophy, this field posits that observed test responses arise from latent psychological constructs [Markus and Borsboom, 2013, Federiakin, 2025]. These constructs may encompass both abilities (e.g., reasoning skills) and personalities (e.g., conscientiousness). The causal framework necessitates rigorous construct definition, requiring traits to be precisely delineated through iterative theory-building and empirical validation. Psychometric test development follows methodical protocols, often structured by frameworks such as Evidence-Centered Design (ECD) [Mislevy et al., 2003]. ECD emphasizes ensuring congruence between test items and theoretical models of the construct, thereby supporting robust inferences about latent traits.
Central to this approach is the prioritization of item quality over quantity. Psychometricians conduct rigorous item analyses to balance precision with practicality, as administering an excessive number of items to participants is often impractical. Advanced statistical models, such as Item Response Theory (IRT) [Embretson and Reise, 2013] and Factor Analysis [Loehlin, 2004], are adopted to estimate latent traits, analyze item performance, and assess model fit. These models require relatively large sample sizes (the number of human participants) to yield stable parameter estimates, as they must disentangle individual differences from measurement error to accurately infer latent traits.
The test results are analyzed to ensure the reliability, validity, predictive power, and explanatory power of the test. Specifically, a well-designed test should: 1) consistently and accurately measure the intended construct, 2) predict performance across a diverse range of related tasks and real-world outcomes, and 3) provide explanatory insight into the observed data. For example, psychometric models often reveal that individual differences across a broad range of cognitive tasks can be captured and explained by a relatively small set of underlying cognitive abilities [Cattell and Horn, 1978].
Benchmark.
In contrast, AI benchmarking is driven by pragmatic goals: evaluating and ranking models based on task performance. Unlike psychometrics, validity is not the primary concern. Instead, benchmarks typically emphasize breadth, scalability, and—especially in the era of foundation models—difficulty. This approach reflects a representativist philosophy, where it is assumed that an extensive set of benchmark items collectively captures all relevant aspects of the abilities demanded by the target task [Federiakin, 2025]. However, constructs like reasoning or knowledge are often ambiguously defined and encompass infinitely many aspects. Benchmarks implicitly operationalize these constructs through ad hoc task selection.
The development of AI benchmarks is usually less labor-intensive, especially when compared with psychometrics on a per-item basis. Test items and their corresponding ground truths are typically drawn from existing datasets, expert curation, or crowdsourced contributions. While this process enables scalability, it risks conflating superficial task performance with deeper cognitive capacities. For instance, a benchmark may assess mathematical reasoning through arithmetic problems without verifying whether models rely on pattern recognition versus symbolic logic [Ahn et al., 2024]. Additionally, LLM benchmarking commonly employs straightforward metrics, such as average accuracy, eschewing the sophisticated latent variable models of psychometrics. This simplicity allows benchmarks to evaluate single models efficiently, bypassing the need for population samples. However, it also limits the depth of insights that can be gleaned from model performance [Federiakin, 2025].
Reliability and stability in benchmarking are primarily achieved through scaling up the test. However, ensuring the quality of each individual item becomes impractical due to the test scale and the rapid pace of model development. For instance, while psychometrics emphasizes the discriminative power of each item, some benchmarks, though initially challenging, are quickly outpaced by continuous model improvements [McIntosh et al., 2024]. Conversely, certain emerging benchmarks are currently too difficult to yield meaningful comparisons [Phan et al., 2025]. Benchmark results are often limited to the specific target task, offering limited generalizability or predictive power across other tasks or real-world applications. These results also pose significant challenges for conducting in-depth, multi-faceted analyses of model capabilities [Wang et al., 2023a].
3.2 Benchmarking with psychometrics-inspired principles
Recent LLM evaluation efforts have drawn inspiration from psychometrics and seek to develop benchmarks that adhere to psychometric principles.
Construct-oriented benchmarking.
Task-oriented benchmarks often entail vast question sets to capture complex abilities. However, in many cases, the benchmarks either fail to fully represent these abilities due to their infinite manifestations or involve extraneous factors that are irrelevant to the target ability [Zhou et al., 2025, Wallach et al., 2025]. Recent research has drawn inspiration from psychometrics and explored the paradigm of construct-oriented evaluation, seeking its discriminative, predictive, and explanatory power. Ilić and Gignac [2024], Federiakin [2025] employ factor analysis to explore the latent variables underlying LLM benchmark performance. Their findings reveal a monolithic factor resembling general intelligence or ability. Federiakin [2025] ranks models based on this discovered factor and highlights its unique advantages over raw benchmark scores. In contrast, similar attempts by Burnell et al. [2023] identify three factors—reasoning, comprehension, and core language modeling—that better explain LLM performance across 27 cognitive tasks. Based on it, Zhu et al. [2024a] integrate the three factors into benchmark items to evaluate multifaceted abilities. This discrepancy between the estimated latent factors in the above findings can be attributed to differences in the models and benchmarks employed [Zhou et al., 2025]. Therefore, rather than relying on statistically derived factors, Zhou et al. [2025] propose a theory-driven hierarchical set of general scales for systematic construct-oriented evaluation. These scales are validated to explain what AI systems can do and predict their performance on novel task instances. Peng et al. [2024] present the Tong Test, a value- and ability-oriented framework, for Artificial General Intelligence (AGI) evaluation. This framework is rooted in dynamic embodied physical and social interactions (DEPSI), and can generate an infinite variety of tasks to evaluate key capabilities including values, learning, and cognition.
Psychometrically rigorous benchmark development.
Beyond defining and analyzing latent constructs, researchers have developed holistic, psychometrically rigorous methods for benchmark development. Liu et al. [2024e] introduce Evidence-Centered Benchmark Design (ECBD), a framework structuring benchmark creation into five modules—capability, content, adaptation, assembly, and evidence—each requiring justification to ensure validity. Through case studies of prominent LLM benchmarks, they demonstrate ECBD’s utility in identifying validity threats. Similarly grounded in psychometrics but with a distinct approach, Fang et al. [2024] propose Psychometrics-Assisted Benchmarking (PATCH), an eight-step process from construct definition to proficiency scoring. When piloted on 8th-grade mathematics, PATCH produced results diverging from traditional benchmarks, offering a more comprehensive evaluation. Building on related principles, Kardanova et al. [2024] adapt Evidence-Centered Design (ECD) to create psychometrically grounded benchmarks. Their application in pedagogy illustrates how this method can reduce data contamination and enhance test interpretability.
Item Response Theory.
Item Response Theory (IRT) is a statistical approach in psychometrics that jointly estimates the latent ability of examinees and the properties (difficulty, discrimination) of test items [Embretson and Reise, 2013]. Translating this framework to LLM evaluation enables researchers to infer latent ability scores, assess item informativeness, and perform more efficient evaluation. Recent research leverages principles from IRT to develop adaptive evaluation frameworks. These methods dynamically calibrate item difficulty based on model performance and weighting items by their inferred difficulty, aiming to achieve accurate evaluations with smaller test sizes and more discriminative items [Lalor et al., 2024, Polo et al., 2024, Zhuang et al., 2023b, a, Guinet et al., 2024]. Building on these approaches, Jiang et al. [2024a], Truong et al. [2025] introduce IRT-based benchmarks that involve learning to estimate item difficulty and learning to generate novel items calibrated to specific difficulty levels. Additionally, research has used IRT-based analyses to explore the alignment between LLM and human response distributions [He-Yueya et al., 2024]. IRT-based evaluations further offer the potential to estimate construct and item parameters on a unified scale, enabling direct comparisons across AI systems and against human norms, even when different test sets are used [Wang et al., 2023a, Fang et al., 2024].
4 Psychometrics for measuring psychological constructs
Figure 2: Examples of psychometric tests for LLMs.
This section delves into the psychological constructs evaluated in LLM Psychometrics. Fig. 2 exemplifies the tests for the involved constructs.
Constructs Theories/inventories Related work
Personality traits Big Five Serapio-García et al. [2023], Jiang et al. [2023], Li et al. [2022b], Zou et al. [2024], Bhandari et al. [2025a], Huang et al. [2024c], Romero et al. [2024], Lee et al. [2024d], Liu et al. [2024a], Karra et al. [2022], Song et al. [2023], Caron and Srivastava [2022], Li et al. [2022a], Kovač et al. [2023], Frisch and Giulianelli [2024], Zheng et al. [2025], La Cava and Tagarelli [2024], Bodroža et al. [2024], Huang et al. [2023a], Gupta et al. [2024], Huang et al. [2023e], Pellert et al. [2024], Ai et al. [2024], Dorner et al. [2023], Klinkert et al. [2024], Petrov et al. [2024], Shu et al. [2024], Ren et al. [2024], Bhandari et al. [2025b], Hilliard et al. [2024], Jiang et al. [2024b], Lu et al. [2023], Sühr et al. [2023], Zhang [2024], Jain et al. [2024]
HEXACO Wang et al. [2025b], Bodroža et al. [2024], Miotto et al. [2022], Barua et al. [2024], Ren et al. [2024], Peereboom et al. [2024]
MBTI Rao et al. [2023], Pan and Zeng [2023], Song et al. [2024b], Cui et al. [2023], La Cava and Tagarelli [2024], Zhang et al. [2024a], Huang et al. [2023c], Ai et al. [2024], Lu et al. [2023], Chen et al. [2024a]
Dark Triad Li et al. [2022b], Romero et al. [2024], Lee et al. [2024d], Li et al. [2022a], Huang et al. [2023e], Barua et al. [2024], Peereboom et al. [2024], Lu et al. [2023]
Other & custom Wen et al. [2024b] (reviews); Mao et al. [2024a], Ai et al. [2024], Jiang et al. [2023], Zeng [2024]
Values Schwartz Pellert et al. [2024], Yao et al. [2025a], Shen et al. [2024], Kovač et al. [2024], Rozen et al. [2024], Zhang et al. [2023a], Yao et al. [2024], Hadar-Shoval et al. [2024], Fischer et al. [2023], Kovač et al. [2023], Miotto et al. [2022], Lee et al. [2024a], Ren et al. [2024], Ye et al. [2025a], Li et al. [2024b], Yao et al. [2025b], Cahyawijaya et al. [2024]
WVS Chiu et al. [2025], Yao et al. [2025a], Kim and Baek [2024], Faulborn et al. [2025]
VSM Kovač et al. [2023], Kharchenko et al. [2024], Zhong et al. [2024], Ren et al. [2024], Ye et al. [2025a]
GLOBE Li et al. [2024d], Karinshak et al. [2024], Ren et al. [2024]
SVO Zhang et al. [2024c]
Other & custom Zhang et al. [2024b], Li et al. [2024d], Biedma et al. [2024], Ye et al. [2025b], Meadows et al. [2024], Jiang et al. [2024a], Moore et al. [2024], Xu et al. [2023], Liu et al. [2024b]
Morality MFT Pellert et al. [2024], Nunes et al. [2024], Abdulhai et al. [2024], Neuman et al. [2025], Ji et al. [2024], Simmons [2023], Tlaie [2024], Aksoy [2024], Yan et al. [2024], Fraser et al. [2022], Zhou et al. [2024a], Yao et al. [2025b], Münker [2024]
ETHICS Jinnai [2024], Rodionov et al. [2023], Karpov et al. [2024], Yu et al. [2023], Albrecht et al. [2022]
DIT Khandelwal et al. [2024], Tanmay et al. [2023]
Other & custom Tanmay et al. [2023], Jin et al. [2022], Huang et al. [2024a], Jin et al. [2024b], Seror [2024], Bonagiri et al. [2024], Garcia et al. [2024], Ahmad and Takemoto [2024], Jiang et al. [2024a], Takemoto [2024], Han [2023], Peterson [2025], Ramezani and Xu [2023], Kucuk and Kocyigit [2023], Neuman et al. [2025], Chiu et al. [2025], Scherrer et al. [2023], Marraffini et al. [2024], Vida et al. [2024], Jinnai [2024], Meijer et al. [2024], Yuan et al. [2024], Liu et al. [2024d], Ohashi et al. [2024], Sachdeva and van Nuenen [2025]
Attitudes & opinions ANES Qi et al. [2024], Jiang et al. [2024d], Yang et al. [2024], Argyle et al. [2023], Bisbee et al. [2024], Sun et al. [2024], Jiang et al. [2022]
ATP Santurkar et al. [2023], Hwang et al. [2023], Tjuatja et al. [2024]
GLES Ma et al. [2024b], Ball et al. [2025], Qi et al. [2024], von der Heyde et al. [2024]
PCT Azzopardi and Moshfeghi [2024], Bernardelle et al. [2024], Wright et al. [2024], Röttger et al. [2024], Hartmann et al. [2023], Rozado [2023], Rutinowski et al. [2024]
Other & custom Ma et al. [2024a] (reviews); Dominguez-Olmedo et al. [2024], Kim and Lee [2023], Sanders et al. [2023], Kalinin [2023], Geng et al. [2024], Xu et al. [2025c], Lee et al. [2024c], Durmus et al. [2023], Rozado [2023], Rosenbusch et al. [2023], Wu et al. [2023], Chalkidis and Brandl [2024], Feng et al. [2023], Ceron et al. [2024], Röttger et al. [2025], Kim et al. [2025b], Faulborn et al. [2025]
Table 2: Representative LLM personality constructs in LLM psychometrics. The main dimensions or focus of the theories/inventories are listed in Table 3.
Theory/inventory What it measures / dimensions
Big Five Five broad personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, Neuroticism
HEXACO Six personality traits: Honesty-Humility, Emotionality, Extraversion, Agreeableness, Conscientiousness, Openness
MBTI Four dichotomies: Extraversion–Introversion, Sensing–Intuition, Thinking–Feeling, Judging–Perceiving
Dark Triad Three negative personality traits: Narcissism, Machiavellianism, Psychopathy
Schwartz Basic human values: 10 or more values (e.g., Self-Direction, Stimulation, Hedonism, Achievement, Power, Security, Conformity, Tradition, Benevolence, Universalism), typically grouped into four higher-order categories (Openness to Change, Self-Enhancement, Conservation, Self-Transcendence)
WVS World Values Survey: Assesses broad cultural values such as traditional vs. secular-rational values, survival vs. self-expression values
VSM Value Survey Module (often Hofstede): Cultural dimensions such as Power Distance, Individualism, Masculinity, Uncertainty Avoidance, Long-Term Orientation, Indulgence
GLOBE Global Leadership and Organizational Behavior Effectiveness: Nine cultural dimensions (e.g., Performance Orientation, Assertiveness, Future Orientation, Humane Orientation, Institutional Collectivism, In-Group Collectivism, Gender Egalitarianism, Power Distance, Uncertainty Avoidance)
SVO Social Value Orientation: Measures individuals’ preferences regarding resource allocation between oneself and others (e.g., prosocial, individualistic, competitive orientations)
MFT Moral Foundations Theory: Five (sometimes six) moral foundations—Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, Sanctity/Degradation, (Liberty/Oppression)
ETHICS Various ethics-related measures assessing moral reasoning, ethical principles, or moral preferences
DIT Defining Issues Test: Assesses moral development and reasoning using moral dilemmas
ANES American National Election Studies: Political attitudes, beliefs, and behaviors in the U.S.
ATP Attitudes Toward Politics: Measures political attitudes and social controversies, often specific to a region or subject
GLES German Longitudinal Election Study: Political attitudes, beliefs, and voting behaviors in Germany
PCT Political Compass Test: Economic (Left/Right) and Social (Authoritarian/Libertarian) political dimensions
Table 3: Personality theories and inventories measured in LLM psychometrics and their main dimensions or focus.
4.1 Measuring personality constructs
LLMs exhibit personality constructs that are not explicitly programmed or trained towards. The emergent constructs critically shape LLM behavior [Hagendorff, 2023], with profound implications for both individuals and broader social groups [Bengio et al., 2024]. Measuring these embedded psychological constructs is essential for understanding behavior, identifying biases, and fostering responsible development.
Table 2 and Table 3 summarize the representative personality constructs that have garnered attention in recent research. Researchers typically select constructs based on their relevance to LLM development and deployment, as well as the applicability of these constructs to AI systems. For instance, Li et al. [2024d] argue that emotional variability in LLMs is not a meaningful construct, given that LLMs lack the biological mechanisms underlying emotions. Conversely, personality and values are considered meaningful for LLMs, as they influence user interactions and model outputs [Serapio-García et al., 2023, Ye et al., 2025a]. This rationale has led to extensive research on personality, values, morality, and attitudes & opinions, as well as some—though less—focus on other constructs such as career selection [Hua et al., 2024], motivation [Chiu et al., 2025, Huang et al., 2023e], and mental health [Reuben et al., 2024, De Duro et al., 2024].
4.1.1 Personality traits
Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life [APA Dictionary of Psychology, ]. In the context of LLMs, personality traits relate to model safety, bias, and toxicity [Zhang et al., 2024a, Wang et al., 2025b], and highly determine user experience [Klinkert et al., 2024].
The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences, including the Big Five [Goldberg, 2013], HEXACO [Ashton and Lee, 2007], Myers-Briggs Type Indicator (MBTI) [Myers et al., 1962], and Dark Triad [Paulhus and Williams, 2002]. When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [Costa and McCrae, 2008], BFI [John et al., 1991], and BFI-2 [Soto and John, 2017], for measuring the Big Five traits; HEXACO-60 [Ashton and Lee, 2009] and HEXACO-100 [Lee and Ashton, 2018] inventories for HEXACO traits; MBTI assessment [Myers, 1985] for MBTI types; and Dark Triad Dirty Dozen scale [Jonason and Webster, 2010] for Dark Triad traits.
Given widespread concerns about the practical relevance of these inventories [Ai et al., 2024], others have adapted them for more real-world scenarios. For example, Bhandari et al. [2025a] contextualize the tests in topic-specific or open-domain conversations; Ai et al. [2024] accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs; Jiang et al. [2023] compile and adapt existing tests into the Machine Personality Inventory; and Mao et al. [2024a] develop the PersonalityEdit inventory using LLMs. However, Peereboom et al. [2024] argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis.
Main findings.
Early models such as GPT-3 score above human averages on Dark Triad scales [Li et al., 2022b, a, Romero et al., 2024]. Even after safety tuning, some models retain certain "dark qualities" [Li et al., 2022b], suggesting that these patterns may be deeply rooted in the pretraining data. More advanced LLMs indicate better level of alignment. They usually demonstrate high Openness, Extraversion, and Agreeableness traits, while low Neuroticism, on the Big Five personality tests [Zou et al., 2024, Bhandari et al., 2025a]. The results align with their design as assistive and helpful entities with emotional stability and engaging personality [Bhandari et al., 2025a, La Cava and Tagarelli, 2024]. Within the MBTI framework, most proprietary LLMs are classified as ENFJ or INFJ types [La Cava and Tagarelli, 2024, Huang et al., 2023c], indicating tendencies toward helpfulness, idealism, and planning capabilities. Personality differences between models are noteworthy. Comparative studies reveal that LLMs from different generations and training methodologies display unique combinations of personality characteristics [Bhandari et al., 2025a, Karra et al., 2022]. LLM personalities also vary across contexts. The same model may display different personality traits across conversational topics [Zou et al., 2024, Song et al., 2023]. This variability challenges the applicability of human personality theories that assume relative trait stability. Multiple studies suggest that prompt design and system instructions substantially influence LLM personality expression [Song et al., 2023, Caron and Srivastava, 2022].
Comparing personality models and suggesting use cases.
The Big Five offers robust empirical foundations and is the most prominent model for personality assessment. However, it is insufficient for predicting comprehensive psychological outcomes [Feher and Vernon, 2021]. HEXACO extends the model by adding Honesty-Humility and provides nuanced insights into moral traits, making it preferable for LLM application contexts demanding integrity [Wang et al., 2025b]. In contrast, MBTI [Myers et al., 1962] categorizes 16 types via dichotomous scales. While popular in career counseling and team-building for its accessibility, its typological approach oversimplifies personality and lacks strong empirical support, limiting its utility in high-stakes decisions [Pittenger, 2005]. The Dark Triad [Paulhus and Williams, 2002] diverges by focusing on maladaptive traits, making it useful in identifying toxic behaviors in LLMs [Barua et al., 2024].
4.1.2 Values
Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group [Schwartz, 1992]. Values offer a powerful lens for understanding LLM behavior. For example, Ye et al. [2025a] show how different values contribute to the safety of LLMs; Liu et al. [2024b] reveal that different spiritual values affect LLMs in social-fairness scenarios; and Sorensen et al. [2024b] demonstrate that standard value alignment reduces distributional pluralism in LLM outputs.
A growing body of research has applied diverse value theories and instruments to assess the value orientations of LLMs. Schwartz’s Value Theory, with its ten basic values and higher-order dimensions, is the most widely used, often via the Schwartz Value Survey (SVS) or Portrait Values Questionnaire (PVQ) [Schwartz, 1992, Schwartz et al., 2001]. Adaptations and contextualized prompts are increasingly common to better suit LLMs’ operational context [Ren et al., 2024, Shen et al., 2024]. The World Values Survey (WVS) has also been used to probe LLMs’ alignment with societal-level value dimensions [Haerpfer et al., 2022, Kim and Baek, 2024]. Cross-cultural frameworks like Hofstede’s Values Survey Module (VSM) and the GLOBE study extend this analysis to workplace and leadership-related cultural dimensions, with both direct and adapted inventories employed [Hofstede, 1984, House et al., 2004, Ren et al., 2024]. Social Value Orientation (SVO) frameworks, using tools like the SVO Slider Measure, focus on LLMs’ prosocial versus proself tendencies [Murphy et al., 2011, Zhang et al., 2024c]. Additionally, researchers have developed localized or custom inventories to capture region-specific or topical values, and some propose novel, bottom-up value taxonomies for LLMs using psycholexical data [Meadows et al., 2024, Xu et al., 2023, Biedma et al., 2024, Ye et al., 2025b]. Collectively, these studies reveal both the adaptability of human value frameworks to LLMs and the need for tailored approaches.
Main findings.
Research indicates that LLMs display distinct and systematic value patterns. According to Schwartz’s Value Theory, LLMs tend to prioritize Self-Transcendence and Conservation. They exhibit stronger inclinations toward Universalism, Benevolence, Conformity, and Security, while opposing Power and Achievement [Rozen et al., 2024, Zhang et al., 2023a, Hadar-Shoval et al., 2024, Fischer et al., 2023]. WVS surveys further suggest that LLMs generally prefer Self-Expression values over Survival values [Chiu et al., 2025]. Studies using the VSM and GLOBE frameworks emphasize LLMs’ strong focus on Humane and Performance Orientation, with moderate Assertiveness [Li et al., 2024b]. Additionally, when assessed through SVO, advanced LLMs predominantly show Prosocial tendencies [Zhang et al., 2024c].
Different models exhibit varied value orientations [Kovač et al., 2024, Chiu et al., 2025, Xu et al., 2023, Li et al., 2024b, Duan et al., 2023]. Versions within the same model family show evolutionary trends in values, potentially influenced by safety alignment, capability advancement, and shifting societal expectations [Duan et al., 2023, Moore et al., 2024, Kim and Baek, 2024]. Generally, larger models align more closely with desirable human values [Shen et al., 2024, Jiang et al., 2024a, Kim and Baek, 2024, Zhong et al., 2024].
Cross-cultural research suggests that LLMs may embody a blend of cultural values, integrating perspectives from diverse backgrounds [Kovač et al., 2023]. However, they generally exhibit a tendency toward Western liberal values [Kim and Baek, 2024]. LLMs can display different values based on profiling prompts [Kharchenko et al., 2024, Zhong et al., 2024, Karinshak et al., 2024]. The context-dependency of their values is concerning and challenges the stability assumptions in human value theories [Shen et al., 2024, Meadows et al., 2024, Moore et al., 2024, Chiu et al., 2025, Xu et al., 2023, Kovač et al., 2023].
Comparing value theories and suggesting use cases.
Schwartz emphasizes universal human motivations; WVS examines societal-level value shifts; VSM and GLOBE address cross-cultural dimensions, with GLOBE emphasizing leadership. SVO, in contrast, focuses on individual-level social preferences, i.e., prosocial vs. proself. While Schwartz and SVO overlap in analyzing individual motivations, the former is broader, and the latter narrowly focuses on decision-making in resource allocation. WVS, VSM, and GLOBE share cultural-level dimensions but differ in focus—WVS on modernization, VSM on workplace-related cultural differences, and GLOBE on leadership and organizational practices. For evaluating individual values in LLMs, use Schwartz for a comprehensive understanding of universal motivations or SVO for interpersonal decision-making studies. In contrast, choose WVS for evaluating broad societal values in LLMs. When evaluating cultural dimensions in LLMs, use VSM for workplace cultural differences or GLOBE for leadership and organizational practices. In addition, Schwartz’s theory bridges individual and societal-level values, making it versatile for mixed-focus studies.
4.1.3 Morality
Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong [Long and Sedley, 1987]. It is crucial to conduct moral assessments of LLMs to ensure their ethical deployment. A large body of research applies the Moral Foundations Theory (MFT) [Graham et al., 2009], primarily through the Moral Foundations Vignettes (MFVs) [Clifford et al., 2015], the Moral Foundations Questionnaire (MFQ) [Graham et al., 2009], the MFQ-2 [Atari et al., 2023], and the Moral Foundations Dictionary (MFD) [Graham et al., 2009]. These papers investigate model bias, alignment with political/moral ideologies, and variation in personal or cultural values, using controlled prompt tests [Tlaie, 2024, Nunes et al., 2024, Abdulhai et al., 2024], persona-driven exploration [Münker, 2024], or evaluating internal moral coherence [Nunes et al., 2024]. Other prominent moral theories and instruments include Kohlberg’s Theory via the Defining Issues Test (DIT) [Kohlberg, 1964], the Consequentialist-Deontological distinction [Beauchamp, 2001], the PEW 2013 Global Attitudes Survey [Center, 2013], and localized moral theories [Liu et al., 2024d, Takeshita et al., 2023]. Specialized datasets are also curated for scalable and comprehensive morality evaluation [Hendrycks et al., 2021, Jinnai, 2024, Marraffini et al., 2024, Jin et al., 2024b]
Main findings.
LLMs are generally characterized by a rationalist and consequentialist focus, often prioritizing harm minimization and fairness [Neuman et al., 2025]. Despite that, they show divergence in ethical reasoning [Neuman et al., 2025] and moral preferences [Tanmay et al., 2023, Jin et al., 2024b, Bonagiri et al., 2024, Meijer et al., 2024]. In some aspects, most LLMs align with human moral standards [Tanmay et al., 2023, Nunes et al., 2024, Takemoto, 2024], which may be attributed to their extensive exposure to conventional ethical values during training. On the other hand, some research presents a less optimistic view, uncovering significant deviations of LLMs from human moral preferences [Marraffini et al., 2024, Vida et al., 2024, Ahmad and Takemoto, 2024]. Regarding the underlying mechanisms of LLMs’ moral reasoning, Nunes et al. [2024], Ji et al. [2024], Simmons [2023] suggest that LLMs primarily exhibit imitation rather than genuine conceptual understanding.
Comparing moral theories/instruments and suggesting use cases.
Instruments based on MFT are well-suited for measuring LLMs’ alignment with intuitive moral principles across cultural or political dimensions. Tools like MFQ, MFQ-2, MFVs, and MFD are particularly effective when investigating ideological leanings, bias, or internal consistency in abstract vs. contextual moral judgments [Nunes et al., 2024]. In contrast, Kohlberg’s DIT explores structured moral reasoning, offering a developmental perspective on how LLMs process ethical dilemmas [Khandelwal et al., 2024, Tanmay et al., 2023]. DIT is valuable for longitudinal studies and evaluating whether models show higher-level moral reasoning (e.g., post-conventional reasoning). Frameworks such as the Consequentialist-Deontological distinction and global surveys like the PEW 2013 study are useful for categorizing ethical tendencies or analyzing cultural homogeneity [Neuman et al., 2025, Meijer et al., 2024]. Localized tools (e.g., CMD, JCM) help assess cultural alignment [Liu et al., 2024d, Takeshita et al., 2023]. Meanwhile, datasets like ETHICS provide morally charged scenarios for detecting normative behavior and fine-tuning LLM responses in high-stakes ethical domains [Hendrycks et al., 2021, Jin et al., 2024b, Marraffini et al., 2024].
4.1.4 Attitudes and opinions
Attitudes are always attitudes about something. This implies three necessary elements: first, there is the object of thought, which is both constructed and evaluated. Second, there are acts of construction and evaluation. Third, there is the agent, who is doing the constructing and evaluating. We can therefore suggest that, at its most general, an attitude is the cognitive construction and affective evaluation of an attitude object by an agent [Bergman, 1998]. We use the term "attitude" to encompass both attitudes and opinions, following [Bergman, 1998, Ma et al., 2024a]. Most research on LLM attitudes examines political attitudes and public opinions [Ma et al., 2024a], which are key cognitive and behavioral foundations in human society and closely tied to model fairness, credibility, and social impact [Durmus et al., 2023, Santurkar et al., 2023, Sanders et al., 2023, Lee et al., 2024c, Hartmann et al., 2023]. For a broader discussion on bias and fairness in LLMs, we refer readers to [Gallegos et al., 2024, Ranjan et al., 2024].
Researchers assess LLMs’ political attitudes using standardized questionnaires and scales from political science and social psychology. US-focused instruments include the American National Election Studies (ANES) [Qi et al., 2024] and the American Trends Panel (ATP) [Santurkar et al., 2023]. For cross-national comparisons, studies use tools like the German Longitudinal Election Study (GLES) [Ma et al., 2024b, Ball et al., 2025, von der Heyde et al., 2024]. The Political Compass Test (PCT) is also widely used to map LLMs within multidimensional political spectrums [Azzopardi and Moshfeghi, 2024, Bernardelle et al., 2024, Hartmann et al., 2023, Röttger et al., 2024].
Other survey instruments include the General Social Survey (GSS) [Kim and Lee, 2023], the American Community Survey (ACS) [Dominguez-Olmedo et al., 2024], the Canadian Election Study (CES) [Sanders et al., 2023], the European Social Survey (ESS) [Geng et al., 2024], the Survey of Russian Elites [Kalinin, 2023], and the Supreme Court Case Political Evaluation (SCOPE) [Xu et al., 2025c]. Additionally, researchers have developed specialized datasets and tailored tools for LLMs, such as OpinionQA [Santurkar et al., 2023], a comprehensive dataset based on ATP surveys; IssueBench [Röttger et al., 2025], a benchmark covering controversial issues with multiple response formats; GlobalOpinionQA [Durmus et al., 2023], which extends political opinion evaluation to cross-cultural contexts; and OpinionGPT [Haller et al., 2023], a web tool that demonstrates how input data biases influence model outputs.
Main findings.
Pretraining data inherently contains socially biased opinions and perspectives, which can amplify political polarization in LLMs [Feng et al., 2023, Xu et al., 2025c]. Most studies identify a misalignment between LLM outputs and human opinions [Santurkar et al., 2023, von der Heyde et al., 2024, Yang et al., 2024, Dormuth et al., 2025], with many concluding that LLMs exhibit a left-leaning political bias [Rozado, 2023, Hartmann et al., 2023, Ceron et al., 2024, Bernardelle et al., 2024, Ma et al., 2024b]. Cross-cultural comparative studies further reveal Western-centric tendencies, demonstrating limited understanding of non-English political perspectives or multi-partisan systems [Qi et al., 2024]. These findings suggest structural limitations in the models’ political understanding. It is also shown that the degree and manifestation of bias vary significantly across contexts and domains. For example, political-electoral propositions exhibit different bias patterns than socioeconomic issues like climate change [Wu et al., 2023]. In contrast, some researchers offer more optimistic perspectives, emphasizing LLMs’ potential to simulate population-level opinions and supplement traditional survey methods [Dominguez-Olmedo et al., 2024, Sanders et al., 2023, Kalinin, 2023, Wu et al., 2023, Jiang et al., 2024d, Argyle et al., 2023, Bisbee et al., 2024, Sun et al., 2024]. They suggest that with appropriate prompt design, calibration methods, and fine-tuning, LLMs can generate opinion distributions closely approximating human group distributions [Jiang et al., 2022, Wu et al., 2023]. Such advancements indicate promising directions for using LLMs in opinion research, though their current limitations and biases must still be acknowledged.
Comparing attitude measurement tools and suggesting use cases.
The ANES offers the most comprehensive framework for American political contexts, making it ideal for studying model performance on U.S. social issues [Jiang et al., 2022]. Cross-national surveys and datasets such as the ESS, GLES, and GlobalOpinionQA provide broader cultural perspectives across diverse sociopolitical environments [Geng et al., 2024, Ball et al., 2025, Ma et al., 2024b, von der Heyde et al., 2024, Durmus et al., 2023]. The ATP is better suited for researching model responses to social controversies [Santurkar et al., 2023], while the PCT serves as a theory-driven political spectrum measurement tool, offering simplified yet intuitive insights into LLM political positioning [Röttger et al., 2024]. Other surveys have distinct regional and sociopolitical focuses: the CES examines political participation in North American multicultural democracies [Sanders et al., 2023]; Survey of Russian Elites provides insights into non-Western political perspectives [Kalinin, 2023]; the SCOPE delivers in-depth assessments of judicial attitudes [Xu et al., 2025c]; and the ACS serves as a tool for researching socioeconomic issues and group representation [Dominguez-Olmedo et al., 2024].
4.2 Measuring cognitive constructs
Traditional NLP benchmarks are insufficient to capture and analyze emergent cognitive constructs [Hagendorff et al., 2024, Ying et al., 2025a]. To address this gap, researchers are adapting related psychometric techniques to evaluate LLM abilities. Hagendorff et al. [2024] introduce the concept of machine psychology, reviewing emergent LLM capabilities and advocating for psychometric evaluation. They categorize these cognitive constructs into four key aspects: heuristics and biases, social interactions, psychology of language, and learning and cognitive capabilities. Structured based on this taxonomy, Table 4 summarizes the representative cognitive constructs measured in recent research.
Aspects Sub-aspects Related work
Heuristics & biases - Yax et al. [2024], Sundaram and Alwar [2024], Momennejad et al. [2023], Echterhoff et al. [2024], Saeedi et al. [2025], Ranaldi and Zanzotto [2024], Macmillan-Scott and Musolesi [2024], Ando et al. [2023], Malberg et al. [2024], Xie et al. [2024b], Hayes et al. [2024], Shah et al. [2024], Chen and Eger [2025], Tang and Kejriwal [2024], Talboy and Fuller [2023], Coda-Forno et al. [2024], Abramski et al. [2023], Binz and Schulz [2023], Xie et al. [2024a], Kumar et al. [2024], Schulze Buschoff et al. [2025], Scholten et al. [2024], Shaikh et al. [2024], Thorstad [2023], Hagendorff et al. [2023], Healey et al. [2024], Bai et al. [2024], Castello et al. [2024]
Social interactions Theory of Mind Sarıtaş et al. [2025], Ma et al. [2023], Mao et al. [2024b] (reviews); Ullman [2023], Sap et al. [2022], Kosinski [2023a], Shapira et al. [2023a], Gandhi et al. [2023], Strachan et al. [2024a], Kosinski [2023b], Moghaddam and Honey [2023], Street et al. [2024], Sclar et al. [2023], van Duijn et al. [2023], He et al. [2023], Holterman and Deemter [2023], Chen et al. [2024d], Zhu et al. [2024c], Hou et al. [2024], Amirizaniani et al. [2024], Soubki et al. [2024], Jones et al. [2024], Lin et al. [2024], Sclar et al. [2024], Sadhu et al. [2024], Lorè et al. [2024], Sarangi et al. [2025], Yu et al. [2025], Tang and Belle [2024], Kim et al. [2025a], Chan et al. [2024], Nickel et al. [2024], Xu et al. [2024a], Leer et al. [2023], Jamali et al. [2023], Li et al. [2023b], Wilf et al. [2023], Chen et al. [2024c], Strachan et al. [2024b], Xu et al. [2025a], Zhang et al. [2025], Shinoda et al. [2025], Jin et al. [2024a], Yang et al. [2025a], Zhou et al. [2023a], Pi et al. [2024], Riemer et al. [2025], Tan et al. [2024].
Emotional Intelligence Sorin et al. [2024], Raj [2024] (reviews); Sabour et al. [2024], Paech [2023], Wang et al. [2023b], Elyoseph et al. [2023], Patel and Fan [2023], Vzorinab et al. [2024], Hu et al. [2025], Li et al. [2024e], Zhao et al. [2024], Chen et al. [2024b], Huang et al. [2024b], Schaaff et al. [2023], Li et al. [2023a], Huang et al. [2023b], Lee et al. [2024f], Welivita and Pu [2024]
Social Intelligence Leng and Yuan [2023], Mittelstädt et al. [2024], Shapira et al. [2023b], Zhou et al. [2024b], Wang et al. [2024d], Liu et al. [2024c], Wang et al. [2024a], Mou et al. [2024], Liang et al. [2024], Zadeh et al. [2019], Xu et al. [2024b], Guo et al. [2023a], Kovač et al. [2024, 2021], Mathur et al. [2024], Schoenegger et al. [2025]
Psychology of language Language comprehension Chang and Bergen [2024] (reviews); Duan et al. [2024a], Wang et al. [2024b], Wilcox et al. [2021], Arehalli et al. [2022], Li et al. [2024a], Steuer et al. [2023], Duan et al. [2024b], Qiu et al. [2024], Bojić et al. [2023], Hu et al. [2023], Ruis et al. [2023], Lee et al. [2024b], Seals and Shalin [2023], Hu and Levy [2023], He et al. [2024], Sun and Wang [2024], Hong et al. [2023], Amouyal et al. [2024], Ide et al. [2024], Lee et al. [2024b], Huff and Ulakçi [2024], Zhou et al. [2023b], Tian et al. [2024]
Language generation Seals and Shalin [2023], Tian et al. [2024], Tang and Kejriwal [2024], Stevenson et al. [2022], Hubert et al. [2023], Chakrabarty et al. [2024], Orwig et al. [2024], Lee and Chung [2024], Boussioux et al. [2024], Miaschi et al. [2024], Bellemare-Pepin et al. [2024], Cai et al. [2024]
Language acquisition Steuer et al. [2023], Shah et al. [2024], Frank [2023b], Riva et al. [2024]
Learning and cognitive capabilities - Niu et al. [2024] (reviews); Galatzer-Levy et al. [2024], Wang et al. [2024e], Dayan et al. [2024], Wu et al. [2025], Sartori and Orrú [2023], Webb et al. [2022], Song et al. [2024a], Zhang et al. [2023b], Wang et al. [2024e], Coda-Forno et al. [2024], Zeng et al. [2024c], Lv et al. [2024]
Table 4: Representative LLM cognitive constructs in LLM psychometrics.
4.2.1 Heuristics and biases
Heuristics and biases are mental shortcuts that simplify decision-making but can introduce systematic errors [Tversky and Kahneman, 1974]. Recent research employs psychometric tools to evaluate rationality and biases in LLM outputs and offer theoretical explanations for these biases.
Early studies, such as Binz and Schulz [2023], found that GPT-3 performs comparably to humans on cognitive ability tests, exhibiting biases like framing and anchoring. More advanced models, such as GPT-4, show fewer System 1 (intuitive) errors and more System 2 (deliberative) reasoning, though biases persist [Hagendorff et al., 2023]. Large-scale evaluations now cover a wide range of biases—including conjunction fallacy, unwarranted beliefs, and loss aversion—using synthetic scenarios and multi-agent dialogue [Malberg et al., 2024, Xie et al., 2024b]. Multimodal models are also being assessed for similar patterns [Schulze Buschoff et al., 2025].
To explain these findings, researchers draw on dual-process theory, linking LLM reasoning to intuitive and deliberative modes [Hagendorff et al., 2023, Yax et al., 2024]. Other frameworks highlight LLMs’ lack of metacognitive abilities, leading to “metacognitive myopia”—the inability to monitor or control reasoning processes, which perpetuates systematic biases [Scholten et al., 2024]. Additional perspectives use cognitive dissonance and elaboration likelihood theories to interpret inconsistencies [Sundaram and Alwar, 2024], while mechanistic approaches leverage influence graphs and Shapley values to trace the origins of biases in model training [Shaikh et al., 2024].
Main findings.
By administering psychometrics to LLMs, studies consistently find that LLMs exhibit cognitive biases superficially similar to humans, such as anchoring, framing, and the conjunction fallacy. Large-scale tests, such as [Echterhoff et al., 2024, Malberg et al., 2024, Xie et al., 2024b], systematically categorize these biases and enable scalable evaluations. Some note that newer, larger, chain-of-thought-enabled models exhibit improved reasoning and bias mitigation [Tang and Kejriwal, 2024, Hagendorff et al., 2023], while others argue that increasing model complexity without deliberate bias mitigation strategies can amplify existing biases [Kumar et al., 2024]. In addition, researchers explore mechanistic interpretability [Shaikh et al., 2024] and present theoretical explanations from cognitive psychology and reasoning theories [Hagendorff et al., 2023, Yax et al., 2024, Scholten et al., 2024]. While LLMs exhibit dual-process reasoning dynamics reminiscent of human cognition [Hagendorff et al., 2023], detailed analysis reveals differences between LLM and human reasoning [Yax et al., 2024].
4.2.2 Social interactions
Researchers apply psychometric tools from social and developmental psychology to assess LLMs’ capabilities in navigating social dynamics. Related evaluation focuses on interconnected dimensions such as Theory of Mind (ToM), Emotional Intelligence (EI), and Social Intelligence (SI).
Theory of Mind (ToM).
ToM is the ability to attribute mental states such as beliefs, intentions, and knowledge to others [Premack and Woodruff, 1978]. Advanced LLMs can simulate ToM-like reasoning under certain conditions, which prompts questions about how these behaviors arise and how robustly they generalize. Foundational studies apply classic psychometrics—false belief tasks—to evaluate ToM in LLMs. [Kosinski, 2023a, b] report that GPT-3.5 and GPT-4 perform at levels matching or exceeding those of children in structured ToM tasks, suggesting the spontaneous emergence of ToM-like behavior as models scale in size. Through mechanistic interpretation, Jamali et al. [2023] offer evidence of parallels between the LLM embeddings and neurons in the human brain.
However, these claims have been challenged. Ullman [2023], Shapira et al. [2023a], Holterman and Deemter [2023] show that minor task changes often cause LLMs’ performance to drop, suggesting reliance on brittle heuristics rather than true semantic understanding. To address this, researchers have developed more robust benchmarks and evaluation protocols featuring procedurally generated, higher-order, or broader ToM tasks [Gandhi et al., 2023, Chen et al., 2024d, He et al., 2023, Xu et al., 2024a, Shinoda et al., 2025]. These benchmarks frequently reveal significant declines in performance on complex tasks, such as 6th-order belief attribution [Street et al., 2024] or faux pas detection [Strachan et al., 2024a].
Some argue these failures stem from a lack of general commonsense reasoning rather than an inability to represent mental states [Pi et al., 2024]. Riemer et al. [2025] further suggest distinguishing between literal and functional ToM, noting that current benchmarks inadequately assess functional ToM. While newer models continue to improve on structured ToM tasks [Strachan et al., 2024a, Street et al., 2024], they remain inconsistent in open-ended, adversarial, or pragmatic reasoning scenarios [Sclar et al., 2024, Yu et al., 2025, Nickel et al., 2024]. Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but still fall short of general human-like social cognition [Strachan et al., 2024a, van Duijn et al., 2023, Jones et al., 2024]. Recent work has also extended ToM evaluation to multimodal settings [Chen et al., 2024c, Strachan et al., 2024b, Jin et al., 2024a] and multi-agent interactions [Li et al., 2023b].
Main findings in ToM.
LLMs exhibit psychometrically measurable ToM-like reasoning, especially when appropriately prompted or structured, but current evidence suggests these capabilities depend on surface-level linguistic cues and lack robustness. The proficiency of LLMs in ToM remains a contentious issue. Interested readers are referred to recent comprehensive reviews on ToM in LLMs [Sarıtaş et al., 2025, Ma et al., 2023, Mao et al., 2024b].
Emotional Intelligence (EI).
EI is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions [Salovey and Mayer, 1990].
Recent work introduces benchmarks to assess EI in LLMs using structured, theory-driven tasks. Tools like EmoBench [Sabour et al., 2024] and EQ-Bench [Paech, 2023] operationalize constructs such as emotional understanding and regulation, often referencing frameworks like Mayer-Salovey-Caruso Emotional Intelligence Test. Other studies adapt established psychometric instruments, including Situational Evaluation of Complex Emotional Understanding (SECEU) [Wang et al., 2023b], the Levels of Emotional Awareness Scale (LEAS) [Elyoseph et al., 2023], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [Patel and Fan, 2023], among others [Huang et al., 2024b]. Advanced models like GPT-4 often match or surpass human baselines in emotional awareness and understanding [Wang et al., 2023b, Elyoseph et al., 2023, Patel and Fan, 2023], though they lack deep reflexive analysis of emotional experience and motivation [Vzorinab et al., 2024]. Recent benchmarks also extend EI evaluation to multimodal settings [Hu et al., 2025].
Main findings in EI.
Advanced LLMs generally perform on par with or better than humans in tasks of EI [Elyoseph et al., 2023, Patel and Fan, 2023], despite evident limitations in several areas, such as displaying artificial or mechanical patterns when expressing empathy [Lee et al., 2024f], the deep reflexive analysis of emotional experiences [Vzorinab et al., 2024], and the misalignment with human emotional behaviors [Huang et al., 2023b].
Social Intelligence (SI).
SI is the ability to understand and manage people [Thorndike and Stein, 1937]. SI in LLMs determines how well these models interpret and respond to social situations. Psychometric tools applied to LLMs yield mixed results. Mittelstädt et al. [2024] use Situational Judgment Tests (SJTs) and find some LLMs outperform humans in expert-rated social appropriateness. However, LLMs still struggle with implicit social understanding: Shapira et al. [2023b] show LLMs perform poorly on faux pas tests, and Xu et al. [2024b] find superficial friendliness often leads to errors in the Situational Evaluation of SI (SESI). The AgentSense benchmark further highlights LLMs’ limitations in complex social interactions, especially regarding high-level needs and private information [Mou et al., 2024].
Recent work expands to multi-agent environments. SOTOPIA [Zhou et al., 2024b] and SOTOPIA-$\pi$ [Wang et al., 2024d] simulate cooperative and competitive social contexts, while the STSS benchmark [Wang et al., 2024a] evaluates SI through task-oriented simulations focused on outcomes and goals. The CogMir framework [Liu et al., 2024c] finds LLMs and humans are similarly consistent in irrational and prosocial decisions under uncertainty. DeSIQ [Guo et al., 2023a] extends SI evaluation to multimodal settings. Critically, Kovač et al. [2024] argue current benchmarks lack developmental grounding and propose the SocialAI school as a more comprehensive framework for studying LLM-based SI.
Main findings in SI.
LLMs show measurable competence in rule-based, socially appropriate behavior, particularly in structured environments. They perform well on: 1) following predefined social norms [Mittelstädt et al., 2024]; 2) completing interactional goals in multi-agent settings [Zhou et al., 2024b, Wang et al., 2024d]; 3) replicating human-like prosocial decisions [Liu et al., 2024c]; and 4) being persuasive, even better than human persuaders [Schoenegger et al., 2025]. Weaker performance is noted in tasks where superficial friendliness causes errors [Xu et al., 2024b], understanding high-level growth needs is essential [Mou et al., 2024], and LLMs must implicitly describe social situations [Shapira et al., 2023b].
4.2.3 Psychology of language
Psycholinguistics, a subfield of psychology, explores how humans comprehend, generate, and acquire language [Carroll, 1986]. Insights from this discipline help evaluate how LLMs process language and mirror human linguistic features.
Language comprehension.
LLM language comprehension is evaluated across multiple linguistic levels—sound, word, syntax, meaning, and discourse. Duan et al. [2024a] introduce a benchmark of 10 psycholinguistic tests to assess these comprehensively. Many studies focus on specific aspects, such as syntactic and semantic processing [Wang et al., 2024b, Arehalli et al., 2022, Wolfman et al., 2024, Wilcox et al., 2021]. Surprisal, a measure of word predictability, is widely used to model processing difficulty in syntactic ambiguities and hierarchical structures [Wilcox et al., 2021, Arehalli et al., 2022, Li et al., 2024a]. Results are mixed: for example, GPT-2-XL shows human-like competence in sound-gender association and implicit causality, but not in sound-shape association [Duan et al., 2024b]. Other evaluation tasks include grammaticality judgment [Qiu et al., 2024, Ide et al., 2024], pragmatic inference [Bojić et al., 2023, Hu et al., 2023, Ruis et al., 2023], argument role processing [Lee et al., 2024b], and discourse comprehension [Duan et al., 2024a]. Advanced models like GPT-4 often match or surpass humans in pragmatic reasoning [Bojić et al., 2023] and grammaticality judgment [Dentella et al., 2024], though performance varies with prompt format [Hu and Levy, 2023] and implicature understanding remains limited [Ruis et al., 2023].
Language generation.
A rich line of research evaluates the creativity of LLM language generation [Tang and Kejriwal, 2024, Stevenson et al., 2022, Hubert et al., 2023, Chakrabarty et al., 2024, Orwig et al., 2024, Lee and Chung, 2024, Boussioux et al., 2024, Bellemare-Pepin et al., 2024], using tests like Guilford’s Alternative Uses Test (AUT) [Guilford et al., 1978] and the Torrance Test of Creative Thinking (TTCT) [Torrance, 1966]. Early models such as GPT-3 lack originality and novelty, while advanced LLMs like GPT-4 surpass the human average in creativity. Notably, Tang and Kejriwal [2024] find LLMs underperform in divergent creativity (e.g., novel uses for familiar objects) but can rival humans in open-ended creative writing. Cai et al. [2024] show that ChatGPT produces human-like responses in most of 12 psycholinguistic tests, but AI-generated analogies often lack human-like linguistic properties [Seals and Shalin, 2023]. LLM-generated stories tend to be positive and lack suspense or narrative diversity compared to human writing [Tian et al., 2024]. Linguistic profiling has also been used to characterize the task-specific language abilities of LLMs [Miaschi et al., 2024].
Language acquisition.
Work on developmental plausibility examines whether LLMs replicate stages of language learning analogous to those in children [Steuer et al., 2023, Shah et al., 2024]. It is shown that regardless of model size, the developmental trajectories of pretrained language models consistently exhibit a window of maximal alignment with human cognitive development [Shah et al., 2024]. Another study by Frank [2023b] draws inspiration from human language development to explain the data inefficiency of LLMs; the relative efficiency of human language acquisition is possibly due to pre-existing conceptual knowledge, multimodal grounding, and the interactive, social nature of their input.
Main findings.
Early foundations that evaluate language models based on BERT and LSTM link computational linguistics and psycholinguistic mechanisms [Ettinger, 2020, Futrell et al., 2019, Arehalli and Linzen, 2020]; these models generally fall short in diverse psycholinguistic tasks. Recent studies extend evaluations to LLMs and broader tasks [Duan et al., 2024a], where advanced LLMs are shown to surpass humans on tasks such as pragmatic reasoning [Bojić et al., 2023] and creative writing [Tang and Kejriwal, 2024]. Some deficiencies still exist, such as limited implicature understanding [Ruis et al., 2023], prompt-sensitive linguistic competence [Hu and Levy, 2023], and lack of human-like psycholinguistic properties [Seals and Shalin, 2023, Tian et al., 2024]. In addition, mixed results are found in the alignment between LLMs and human linguistic cognition [Duan et al., 2024b, Wolfman et al., 2024] and language acquisition [Shah et al., 2024, Frank, 2023b].
4.2.4 Learning and cognitive capabilities
Psychometrics of learning and cognitive capabilities measures mental functions such as memory, reasoning, problem-solving, and comprehension to identify cognitive strengths and weaknesses. Recent studies adapt human psychometric tools to evaluate LLMs’ cognitive abilities. For example, Galatzer-Levy et al. [2024] use the Wechsler Adult Intelligence Scale (WAIS-IV) to show LLMs generally reach top human levels in verbal comprehension and working memory, though multimodal models lag in visual reasoning. Raven’s Progressive Matrices and other fluid intelligence tests reveal LLMs can match or surpass humans in analogical reasoning [Sartori and Orrú, 2023, Webb et al., 2022].
However, limitations persist. The Montreal Cognitive Assessment (MoCA) shows most LLMs display mild cognitive impairment, especially in visuospatial and executive tasks [Dayan et al., 2024]. Abstraction and Reasoning Challenge (ARC) further highlights deficits in abstract problem-solving [Wu et al., 2025]. Some works question the validity of current tests [Löhn et al., 2024, Li et al., 2024d, Zhang et al., 2023b], prompting new benchmarks and methodologies [Coda-Forno et al., 2024, Wang et al., 2024e, Zhuang et al., 2023a, Zeng et al., 2024c, Song et al., 2024a].
Developmental perspectives also emerge: Wang et al. [2024e] leverage Piaget’s Theory of Cognitive Development to track human-like progression in LLMs, finding advanced models (e.g., GPT-4) reach cognitive abilities comparable to 20-year-old humans. Shah et al. [2024] identify developmental profiles in LLMs paralleling human development via psychometric theories like Cattell-Horn-Carroll.
Main findings.
LLMs demonstrate strong performance on verbal comprehension, working memory, and analogical reasoning, often reaching or surpassing high human percentiles when evaluated with adapted psychometric tools like WAIS-IV and Raven’s Matrices [Galatzer-Levy et al., 2024, Webb et al., 2022]. However, they exhibit notable cognitive deficits, particularly on benchmarks like MoCA and ARC tasks [Wu et al., 2025, Dayan et al., 2024]. Several works introduce developmental and cognitive benchmarks (e.g., CogLM, CogBench) grounded in psychological theories to track reasoning and adaptability [Wang et al., 2024e, Coda-Forno et al., 2024]. While emergent reasoning appears at scale, challenges remain in interpretability, test validity, and generalization beyond surface-level statistical patterns [Li et al., 2024d, Löhn et al., 2024].
5 Psychometric evaluation methodology
Figure 3: Overview of LLM psychometric evaluation methodologies.
This section examines the methodologies employed in LLM Psychometrics. As illustrated in Fig. 3, the methodological framework encompasses four key components: test formats (§ 5.1), data and task sources (§ 5.2), prompting strategies (§ 5.3), and model output and scoring (§ 5.4). The configuration of inference parameters (§ 5.5) is additionally discussed.
5.1 Test format
Table 5: Overview of test formats: structured tests, open-ended conversations, and agentic simulations. For each format, representative examples are provided for each construct.
Construct Structured tests Open-ended conversations Agentic simulations
Personality traits Jiang et al. [2023], Serapio-García et al. [2023] Jiang et al. [2024b], Zheng et al. [2025] Frisch and Giulianelli [2024], Huang et al. [2024c]
Values Kovač et al. [2023], Zhong et al. [2024] Ren et al. [2024], Yao et al. [2025b] Chiu et al. [2025], Shen et al. [2024]
Morality Ji et al. [2024], Abdulhai et al. [2024] Sachdeva and van Nuenen [2025], Neuman et al. [2025] Chiu et al. [2025], Nunes et al. [2024]
Attitudes & opinions Bernardelle et al. [2024], Argyle et al. [2023] Röttger et al. [2024], Wright et al. [2024] -
Heuristics and biases Hagendorff et al. [2023], Yax et al. [2024] Healey et al. [2024], Wen et al. [2024a] Bai et al. [2024], Xie et al. [2024b]
Social interaction Sabour et al. [2024], Mittelstädt et al. [2024] Welivita and Pu [2024], Elyoseph et al. [2023] Zhou et al. [2024b], Wang et al. [2024d]
Psychology of language Duan et al. [2024a], Amouyal et al. [2024] - -
Learning and cognitive capabilities Wu et al. [2025], Coda-Forno et al. [2024] Galatzer-Levy et al. [2024] Lv et al. [2024]
Test formats of LLM psychometric evaluation can be categorized into structured tests, open-ended conversations, and agentic simulations. Table 5 collects exemplary works illustrating the test formats used for each construct.
5.1.1 Structured tests
Structured tests feature predefined instructions, questions, and response formats. These items may include alternative-choice questions, multiple-choice questions, rating scales, and short-answer questions. When adapting structured psychometric tests for LLMs, it is common practice to retain the original items and simply reformat them as prompts.
For example, personality and values are often measured with Likert scales (e.g., BFI statements like “I see myself as someone who is outgoing, sociable”), while morality is assessed using tools like the DIT or MFQ, which require decisions and agreement ratings. Heuristics and biases are often tested with scenario-based choices. Social interaction and psycholinguistic abilities are evaluated with multiple-choice, forced-choice, or masked word prediction. Cognitive tests also involve short-answer questions (e.g., WAIS-IV Digit Span).
Structured psychometric tests are advantageous due to their scalability, objectivity, and automated scoring. However, critical perspectives underline their limitations, such as gaps in real-world applicability; biases; data contamination; and issues with reliability, validity, and depth of insights.
5.1.2 Unstructured tests
Unstructured tests probe LLMs’ personalities and abilities by analyzing their free-form responses, usually with rationale and justification, to user queries, or by contextualizing LLMs to observe their decision-making in real-world scenarios.
Open-ended conversations.
Unstructured testing often engages LLMs in open-ended, typically single-turn conversations that elicit specific constructs, mirroring real-world human-LLM interactions. For example, Jiang et al. [2024b] have LLMs generate long-form narratives to reveal personality traits. ValueBench uses advice-seeking queries to assess LLMs’ influence on users’ values [Ren et al., 2024], while Sachdeva and van Nuenen [2025] present moral dilemmas from Reddit, prompting LLMs to provide community-style moral judgments. To probe political attitudes, Röttger et al. [2024] convert PCT multiple-choice items into forced-choice and open-ended formats. For cognitive constructs, Healey et al. [2024] design pipelines to detect nuanced biases in free responses, and Welivita and Pu [2024] evaluate empathy by asking LLMs to respond to emotional scenarios. Many cognitive tests, such as WAIS-IV vocabulary and comprehension, also use open-ended questions [Galatzer-Levy et al., 2024].
Agentic simulations.
Advanced unstructured tests evaluate LLMs as agents in complex, contextualized role-playing scenarios, analyzing their decision-making in dynamic environments. Huang et al. [2024c] design LLM-Agents with distinct personalities and analyze them by replicating human-like behaviors in agentic simulations. Shen et al. [2024] measure values using four real-world vignettes: collaborative writing, education, public sectors, and healthcare. Chiu et al. [2025] present value and moral dilemmas to LLMs, examining their decisions and rationales. Cognitive construct tests are often more complex, extending to multi-agent settings; for example, Xie et al. [2024b], Bai et al. [2024] identify cognitive biases in multi-agent communication. SOTOPIA simulates open-ended social interactions among LLM agents to evaluate social intelligence [Zhou et al., 2024b, Wang et al., 2024d]. Lv et al. [2024] benchmark cognitive dynamics by having LLMs repeatedly complete cognitive questionnaires and reason over information flows.
Unstructured tests assess LLMs in free-form, contextualized scenarios, offering greater ecological validity and revealing complex behaviors—such as nuanced reasoning, subtle biases, and dynamic social interactions—not captured by structured formats. However, they pose challenges in standardization, scoring, and reproducibility.
5.2 Data and task sources
Table 6: Overview of data and task sources: established inventories, custom-curated items, and synthetic items. For each data source, representative examples are provided for each construct.
Construct Established inventories Custom-curated items Synthetic items
Personality traits Serapio-García et al. [2023], Barua et al. [2024] Ai et al. [2024], Jiang et al. [2023] Zeng [2024]
Values Fischer et al. [2023], Miotto et al. [2022] Shen et al. [2024], Meadows et al. [2024] Ye et al. [2025a], Moore et al. [2024]
Morality Abdulhai et al. [2024], Khandelwal et al. [2024] Hendrycks et al. [2021], Jin et al. [2022] Scherrer et al. [2023], Liu et al. [2024d]
Attitudes & opinions Argyle et al. [2023], Bernardelle et al. [2024] Durmus et al. [2023], Ceron et al. [2024] Wan and Chang [2024]
Heuristics and biases Hagendorff et al. [2023], Binz and Schulz [2023] Momennejad et al. [2023], Ando et al. [2023] Malberg et al. [2024], Xie et al. [2024b]
Social interaction [Ullman, 2023, Chen et al., 2024d] Shapira et al. [2023a], Street et al. [2024] Gandhi et al. [2023]
Psychology of language Cai et al. [2024], Duan et al. [2024a] He et al. [2024], Pérez-Mayos et al. [2021] -
Learning and cognitive capabilities Coda-Forno et al. [2024], Dayan et al. [2024] Song et al. [2024a], Wu et al. [2025] -
Data and task sources LLM Psychometrics can be 1) drawn from established psychometric inventories, 2) human-authored and custom-curated, and 3) synthesized by AI models. Table 6 provides an overview of the data and task sources for each construct.
Established inventories.
Established psychometric inventories offer standardized, well-validated tools for LLM evaluation. Common examples include BFI and HEXACO for personality [Serapio-García et al., 2023, Barua et al., 2024]; SVS, PVQ, WVS, and VSM for values [Fischer et al., 2023, Miotto et al., 2022]; MFT and DIT for morality [Abdulhai et al., 2024, Khandelwal et al., 2024]; ANES and PCT for political attitudes [Argyle et al., 2023, Bernardelle et al., 2024]; CRT for heuristics and biases [Hagendorff et al., 2023, Binz and Schulz, 2023]; and False-Belief Tasks for ToM [Ullman, 2023, Chen et al., 2024d]. While these inventories enable straightforward, standardized assessment, their use with LLMs raises concerns about data contamination, response bias, and limited ecological validity for complex AI tasks [Ye et al., 2025a].
Custom-curated items.
Human-authored, custom-curated items offer tailored psychometric tests that are often more relevant to LLMs, enabling exploration of novel capability dimensions. For example, Shen et al. [2024] find SVS and PVQ insufficient for LLM value alignment, so they create 11 AI-informed value statements based on a systematic review of alignment literature. Ceron et al. [2024] highlight the unreliability of standard political questionnaires for LLMs and introduce tests using annotated voting-advice questionnaires from seven EU countries to assess stance consistency. For morality, Hendrycks et al. [2021] curate a dataset of over 130,000 open-world scenarios requiring moral judgments. In ToM, Soubki et al. [2024] develop a dataset from real spoken dialogues to address gaps between synthetic benchmarks and human behavior. While custom-curated items offer high ecological validity, their development and validation are labor-intensive, limiting scalability and diversity.
Synthetic items.
Synthetic items—generated primarily by LLMs—are an emerging approach in LLM Psychometrics, enabling large-scale, diverse, and contextually rich test content. This method requires careful prompt engineering and validation. Some studies adapt established inventories to improve ecological validity, reduce data contamination, and enhance scalability. For instance, Ren et al. [2024] convert self-report items into advice-seeking queries to better reflect real-world human-AI interactions. Bhandari et al. [2025b] rewrite established items into semantically equivalent forms, validating them via semantic embedding similarity thresholds. Zhu et al. [2024a] diversify cognitive test items to enable more comprehensive analysis of LLM cognitive abilities. Other work generates synthetic tests entirely from scratch. Hadar-Shoval et al. [2025] demonstrate that LLM-generated cognitive tests correlate strongly with human performance on established tests. Ye et al. [2025a] create value-eliciting prompts and confirm the reliability and validity of the resulting measurements. Jiang et al. [2024a] introduce generative evolving testing, where an LLM-based Item Generator produces items of specified difficulty. Chiu et al. [2025] use GPT-4 to generate moral dilemmas for evaluating LLM moral judgment. For social intelligence, agentic simulations prompt LLMs to generate components of social interactions [Zhou et al., 2024b, Mou et al., 2024].
5.3 Prompting strategies
Many structured tests employ standard test prompts, reformatted from those used for human participants. Others involve various prompting strategies.
Role-playing prompts.
Role-playing prompts (persona or profiling prompts) embed demographic or personal attributes into the LLM context during testing. These prompts enable the generation of multiple simulated participants from a single LLM for statistical analysis. For example, Serapio-García et al. [2023] prepend persona instructions to each item to elicit diverse responses, allowing examining reliability and validity within one LLM. Similarly, Ye et al. [2025b] use value-anchoring prompts to create hundreds of LLM-based participants for value measurement, supporting the development of an LLM-specific value system. LLMs can represent a superposition of personalities, values, and perspectives [Kovač et al., 2023], prompting research into their adaptability. Role-playing prompts show that LLMs can shift between personality types, though steerability varies by model and dimension [Lu et al., 2023, Jiang et al., 2023, La Cava and Tagarelli, 2024, Li et al., 2024d]. Such prompts are also used to test the stability of LLM values [Kovač et al., 2024, Rozen et al., 2024] and to measure how well LLMs can mimic moral or political profiles and their susceptibility to related biases [Münker, 2024, Simmons, 2023, Wright et al., 2024]. Cognitive benchmarks, especially for social interactions, often use persona prompts in agentic simulations [Zhou et al., 2024b, Mou et al., 2024, Huang et al., 2023b], with task performance depending on the LLM’s ability to embody these roles. These prompts influence social-cognitive reasoning when controlling for other variables [Tan et al., 2024]. While role-playing capabilities enable diverse perspectives useful for social simulations, they can also lead to unstable values and perspectives, potentially compromising test validity and complicating value alignment.
Performance-enhancing prompts.
Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations, as recommended by Hagendorff [2023]. Chain of Thought (CoT) prompting [Wei et al., 2022] reduces biases [Hagendorff et al., 2023], enhances social intelligence [Shapira et al., 2023a], and improves cognitive capabilities [Coda-Forno et al., 2024]. Its variant Emotional CoT [Li et al., 2024e] boosts EI, while emotional prompts [Li et al., 2023a] enhance general cognitive abilities through emotional stimuli. Few-shot prompting [Brown et al., 2020] has demonstrated improvements in ToM performance [Moghaddam and Honey, 2023]. Research has also developed construct-specific prompting strategies. Zhou et al. [2024a] propose MFT-guided reasoning prompts for moral alignment, while Echterhoff et al. [2024], Sumita et al. [2024] focus on self-debiasing approaches. ToM-specific enhancements include SymbolicToM [Sclar et al., 2023] and temporal decomposition-based prompting [Sarangi et al., 2025, Hou et al., 2024]. Additionally, Zhao et al. [2025a] develop a framework using self-reflection prompts to examine both explicit and implicit social biases, where explicit bias measurement reflects implicit bias.
Prompt perturbation and adversarial attacks.
Prompt perturbations can test the robustness of LLMs’ personality traits and cognitive abilities. Researchers have investigated whether LLMs can maintain stable personality, values, opinions, and cognitive abilities when item options are reordered [Schelb et al., 2025, Lee et al., 2024d], prompts are rephrased [Lee et al., 2024d, Fraser et al., 2022, Strachan et al., 2024a], prompt formats are altered [Moore et al., 2024, Schelb et al., 2025], or different languages are used [Moore et al., 2024, Cahyawijaya et al., 2024]. For instance, Faulborn et al. [2025] introduce 30 variations of prompts to evaluate the political biases of LLMs. In addition, Wen et al. [2024a] propose psychometric-inspired adversarial attacks to uncover implicit biases in LLMs. Similarly, Li et al. [2024d] utilize persuasive adversarial prompts (PAP) [Zeng et al., 2024a] to test the robustness of LLMs’ values. Since LLMs are sensitive to these perturbations, researchers have begun to scrutinize the reliability of the test results obtained under standard conditions [Ye et al., 2025a, Röttger et al., 2024, Dominguez-Olmedo et al., 2024, Strachan et al., 2024a].
5.4 Model output and scoring
5.4.1 Closed-ended output and scoring
LLM performance on structured tests is typically evaluated using logit-based probabilistic analysis and closed-ended output scoring. In multiple-choice or Likert-scale settings, some studies analyze token-level logits—especially the first token—to infer latent traits, detect response entropy, or compare LLM output distributions against human data [Pellert et al., 2024, Santurkar et al., 2023, Dominguez-Olmedo et al., 2024, Arehalli et al., 2022]. Psycholinguistic tests often use logit-based metrics such as surprisal (the negative log probability of a token), which quantifies predictability and cognitive effort, allowing researchers to align LLM uncertainty patterns with those of humans [Wang et al., 2024b, Steuer et al., 2023].
Conversely, closed-ended outputs, which are explicit numerical scores or categorical selections, can be analyzed using predefined scoring protocols. For Likert-scale responses, scores are typically averaged or aggregated based on established rubrics (e.g., VSM13 [Ye et al., 2025a] and DIT [Khandelwal et al., 2024]). In standardized cognitive tests, model outputs are evaluated against ground-truth labels (e.g., accuracy in arithmetic tasks) or rule-based criteria (e.g., established standards in the Verbal Comprehension Index) [Galatzer-Levy et al., 2024].
5.4.2 Open-ended output and scoring
Scoring open-ended outputs is more complex and typically falls into three categories: 1) rule-based, 2) model-based, and 3) human scoring.
Rule-based scoring.
Rule-based scoring predominantly relies on the lexical hypothesis [Allport and Odbert, 1936], analyzing the presence and frequency of specific keywords or phrases. For example, Jiang et al. [2024b] apply LIWC (Linguistic Inquiry and Word Count) features [Pennebaker et al., 2001] to evaluate LLM personalities, and Fischer et al. [2023] employ a theory-driven value dictionary [Ponizovskiy et al., 2020] for value measurement. However, such lexicon-based approaches struggle to capture semantic nuance [Ye et al., 2025a]. Some rule-based methods go beyond keyword matching; for instance, Healey et al. [2024] detect bias by checking for deviations from equivalent treatment in responses.
Model-based scoring.
Model-based scoring is prevalent in unstructured testing for its flexibility and scalability. Many studies train models to evaluate LLM responses, such as fine-tuning BERT variants on the MyPersonality dataset for personality scoring [Hilliard et al., 2024]. Others fine-tune LLMs with psychometric inventories or human annotations to classify value valence [Ye et al., 2025a, Yao et al., 2024, 2025b, Sorensen et al., 2024a]. Scoring can be conducted at the item level (e.g., Generative Psychometrics: item parsing, scoring, aggregation [Ye et al., 2025a]) or at the response level, evaluating the entire output [Yao et al., 2024, 2025b]. Some models are tailored to specific frameworks (e.g., Schwartz’s Value Theory [Yao et al., 2024]), while others are trained to be generalists by leveraging broader fine-tuning data and LLM prior knowledge [Ye et al., 2025a]. Model-based evaluators are especially common for personality and value assessment, where textual cues map readily to target dimensions. Many studies now use the LLM-as-a-judge approach [Gu et al., 2024] for scoring open-ended responses. For instance, Li et al. [2024d], Zheng et al. [2025] prompt LLMs to rate responses on personality, ToM, and motivation, validating consistency with other LLMs or human raters. This method is especially useful in unstructured simulation settings, such as SI measurement across goal completion, relationship maintenance, and social rule adherence [Zhou et al., 2024b, Wang et al., 2024d, a, Mou et al., 2024]. Embedding models are also used to measure similarity between LLM outputs and prototypical examples [Huang et al., 2024c, Cahyawijaya et al., 2024, Amirizaniani et al., 2024].
Human scoring.
Human scoring is employed when rigorous evaluation, adhering strictly to standard psychometric manuals, is necessary. For instance, Elyoseph et al. [2023] engage psychologists to assess the contextual suitability of LLM responses using the Levels of Emotional Awareness Scale (LEAS). Castello et al. [2024] conduct an examination of cognitive biases in LLMs through human evaluation and linguistic comparison. Healey et al. [2024] present a semi-automated pipeline to classify LLM responses into nuanced bias types. Other examples include ToM evaluation in open-ended responses [Amirizaniani et al., 2024] and cognitive evaluation based on the Montreal Cognitive Assessment (MoCA) [Dayan et al., 2024].
5.5 Inference parameters
Evaluation and validation depend on LLM inference parameters. Greedy decoding yields deterministic, reproducible outputs but limits diversity, while sampling-based decoding (adjusting temperature, top-k, top-p) introduces stochasticity and diversifies responses. Sampling-based decoding may reveal a broader spectrum of latent traits, opinions, or cognitive strategies, but also introduce variability that complicates psychometric validation. Conversely, deterministic settings enhance reproducibility but may obscure the model’s full range of capabilities or biases. Most studies report and control inference settings for fair comparisons, and some analyze sensitivity to these parameters. Transparent reporting and careful parameter selection are essential; both deterministic and stochastic settings should be considered to fully understand the implications of test results.
6 Psychometric validation
Figure 4: Overview of psychometric validation: reliability and consistency, validity, and standards and recommendations.
Unlike AI benchmarking, which focuses on system performance, psychometrics prioritizes theoretical grounding, standardized protocols, and reproducibility. Psychometric validation ensures that tests are reliable, valid, and fair. As an emerging field, LLM Psychometrics lacks standardization in test design and administration, and recent work is beginning to address these issues. We outline the landscape of psychometric validation in Fig. 4.
6.1 Reliability and consistency
Reliability, a core aspect of psychometric validation, measures how consistently a test performs—over time (test-retest), across versions (parallel forms), and among evaluators (inter-rater). Researchers have adapted these metrics for LLMs. For example, Li et al. [2024d] propose a benchmark covering five reliability forms: internal consistency, parallel forms, inter-rater, option position robustness, and adversarial attack robustness. Other studies [Zheng et al., 2025, Huang et al., 2023a, Ceron et al., 2024, Shu et al., 2024] examine LLM output consistency across repeated trials, prompt variations, and languages. Notably, some metrics capture both random and systematic errors, so inconsistencies can undermine both reliability and validity.
Some advanced LLMs show stable Big Five personality trait responses across settings [Huang et al., 2023a], with larger and instruction-tuned models achieving even higher reliability using role-playing prompts [Serapio-García et al., 2023]. LLMs also display consistent value orientations across paraphrased, related, and translated questions, as well as different response formats [Moore et al., 2024]. Strong inter-rater reliability is reported when LLMs act as judges, though personality stability and value consistency vary by model and cultural context [Li et al., 2024d]. However, other studies report less favorable results. LLMs are notoriously sensitive to prompt variations: small changes in prompt form [Röttger et al., 2024, Ren et al., 2024], option order [Li et al., 2024d, Dominguez-Olmedo et al., 2024], or syntax [Hu and Levy, 2023] can systematically shift outputs, undermining parallel forms reliability. Studies also find inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [Hu and Levy, 2023], and between forced-choice and free-form responses in value [Kovač et al., 2024, Ye et al., 2025a] and political opinion surveys [Röttger et al., 2024]. In cognitive tests, minor changes to ToM tasks often cause poor performance, indicating low parallel forms reliability and reliance on fragile heuristics [Ullman, 2023, Shapira et al., 2023a, Holterman and Deemter, 2023]. The Words and Deeds Consistency Test (WDCT) [Xu et al., 2025b] further shows LLMs’ verbal and behavioral responses are inconsistent across domains.
Reliability findings are mixed and depend on factors such as construct, reliability type, test format, prompting, and model. LLMs are generally more reliable in personality assessments [Huang et al., 2023a, Serapio-García et al., 2023] than in political opinion evaluations [Dominguez-Olmedo et al., 2024, Röttger et al., 2024]. Inter-rater agreement varies across instruments and domains [Bodroža et al., 2024]. LLMs often show high internal consistency (Cronbach’s alpha $>0.8$) in closed-choice personality tests [Huang et al., 2023a, Serapio-García et al., 2023, Zheng et al., 2025], but low parallel form reliability due to prompt sensitivity [Li et al., 2024d, Shu et al., 2024, Gupta et al., 2024]. Consistency improves in structured tests with strong prompt control [Rozen et al., 2024, Serapio-García et al., 2023, Klinkert et al., 2024], but declines with less standardized prompts [Petrov et al., 2024] or controversial topics [Moore et al., 2024]. Reliability also varies by model: base models are more consistent than fine-tuned ones on value-laden questions [Moore et al., 2024], and higher model safety correlates with greater value consistency [Ye et al., 2025b]. Advanced models also perform better in role-playing personality tests, increasing reliability [Serapio-García et al., 2023].
6.2 Validity
Validity assesses whether a test truly measures its intended construct. Recent work in LLM Psychometrics has explored various facets of validity, highlighting both challenges and emerging solutions.
6.2.1 Content validity
Content validity requires that a test fully represents the intended construct. Key challenges include data contamination from established tests and insufficiently vetted novel items. Directly applying human tests to LLMs risks contamination, as LLMs may have seen test items or similar content during training [Hagendorff et al., 2024, Jiang et al., 2024c]. This can inflate performance or bias trait measurement. To mitigate this, researchers reformat tests or generate new items; see § 5.2. On the other hand, reformatted or newly generated items may inadequately capture the construct or introduce extraneous factors. This may be due to a less rigorous crowd-sourcing process for benchmark curation, or the inherent biases and limited capabilities of the AI models used for stimulus synthesis. Evaluating content validity for custom-curated and model-generated items is crucial but rarely conducted in LLM Psychometrics.
6.2.2 Construct validity
Construct validity determines if a test truly measures its intended construct. Alaa et al. [2025] highlight its critical yet often lacking role in medical LLM benchmarks. Broader literature suggests that construct validity is challenged by LLMs’ unique internal abstractions, systematic response patterns, and social desirability bias.
Unique internal abstractions
LLMs and humans may fundamentally differ in their internal representations of psychological constructs. Kovač et al. [2023] argue that LLMs display contextually adaptive traits, unlike the stable characteristics seen in humans. Sühr et al. [2023] show that LLM personality types do not align with the clear separation of Big Five traits found in humans. Similarly, Peereboom et al. [2024] find that human-oriented HEXACO tests are inapplicable to LLMs, as key personality factors may be absent. For values, Ye et al. [2025a] report that LLM self-reports do not fit Schwartz’s model, and while open-ended responses improve alignment, it remains an imperfect fit; Hadar-Shoval et al. [2024] also find LLMs encode values distinct from humans. Overall, researchers agree that robust LLM psychometric evaluation requires new operational definitions, distinct from or only loosely analogous to human traits.
In this regard, pioneering research presents value systems tailored to LLMs’ unique internal abstractions [Ye et al., 2025b, Biedma et al., 2024]. Ye et al. [2025b] introduce a generative psycho-lexical approach based on GPV value measurement [Ye et al., 2025a], demonstrating through various tasks that their value factors are valid and outperform Schwartz’s system. The distinct nature of LLM constructs requires rigorously designed tests: Ma et al. [2025] present the Core Sentiment Inventory (CSI) for measuring LLM emotional tendencies and personality; Fang et al. [2024] create a psychologically grounded mathematics benchmark to improve validity, model item difficulty, and enable human comparison; and Lee et al. [2024d] propose the TRAIT test, achieving higher validity and reliability than other personality measures.
Response set.
Response set refers to systematic answer patterns regardless of item content. Some LLMs show option position bias, favoring responses labeled "A" in political opinion tests [Dominguez-Olmedo et al., 2024, Li et al., 2024d]. After correcting for this bias, Dominguez-Olmedo et al. [2024] find LLM responses to political questions are uniformly random, rendering demographic alignment analyses meaningless. Similar findings are observed in LLM personality [Song et al., 2023, Sühr et al., 2023] and value tests [Ye et al., 2025a]. Sühr et al. [2023] also note LLMs often affirm both the original and reverse-coded items, while Ye et al. [2025a] observe some LLMs consistently prefer low or high ratings. These observations suggest self-report tests may be unsuitable for LLM Psychometrics.
Social desirability bias.
Social desirability bias—the tendency to give socially acceptable rather than genuine responses—is observed in LLMs as in humans. Salecha et al. [2024] find LLMs skew toward favorable traits in personality tests, inflating extraversion and lowering neuroticism scores. Similarly, Ye et al. [2025b] show LLMs avoid reporting less desirable values (e.g., hedonism) in direct self-reports [Biedma et al., 2024], but reveal them in indirect, contextual prompts. While tuning for social desirability is important, overemphasis risks producing homogeneous models that lack diverse perspectives and fail to meet varied user needs.
Cross-lingual tests.
Cross-lingual psychometric tests reveal notable inconsistencies in LLM measurement results across languages [Romero et al., 2024]. Despite multilingual capabilities, LLMs often lack consistent psychological trait assessments between languages [Cahyawijaya et al., 2024]. These inconsistencies complicate validation, requiring careful separation of model biases, translation issues, and cultural differences represented by the models.
6.2.3 Criterion and ecological validity
Criterion validity examines how well test results correspond to external standards, while ecological validity concerns their relevance to real-world scenarios. In LLM Psychometrics, external standards often coincide with real-world evaluation results. Studies find that value orientations from forced-choice tests frequently diverge from those observed in human-LLM interactions [Ren et al., 2024, Ye et al., 2025a], with similar gaps in personality [Ai et al., 2024] and morality [Nunes et al., 2024]. As a result, researchers advocate linking LLM personality to safety [Zhang et al., 2024a], measuring values within real-world interactions [Ren et al., 2024], and using open-ended dialog to improve construct validity [Ye et al., 2025a]. Further work connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards [Ye et al., 2025b].
6.3 Standards and recommendations
To address key challenges, researchers have proposed standards and recommendations to guide LLM Psychometrics and establish methodological rigor. Frank [2023a] suggests leveraging developmental psychology to understand LLM cognition, recommending simplified and novel stimuli to prevent training data contamination. Löhn et al. [2024] critique current practices and outline seven requirements for valid assessment: reliability, validity, suitability to LLM capabilities, non-disclosure (to avoid contamination), and fairness—including valid comparisons, translations, and transparency. Their review of 25 studies reveals these principles are often overlooked.
More recently, Hagendorff et al. [2024] advocate for procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options, and diverse scoring methods to reduce contamination and improve reliability. They also recommend deterministic settings for reproducibility, automated evaluation tools, manual review of unreliable outputs, and statistical analysis of results. Vaugrante et al. [2024] highlight poor replicability in recent work and offer four main recommendations: (1) ensure benchmark validity, sufficient task coverage, standardization, prompt control, and alignment with research goals; (2) use standardized methods, avoid cherry-picking, maintain statistical transparency, document setups, and define consistent metrics; (3) monitor model changes, account for model diversity, adjust benchmark difficulty as models advance, and document model versions and experiment dates; (4) standardize scoring and verification, with clear guidelines and rubrics for implementers. Schelb et al. [2025] introduce a framework for robust, flexible, and reproducible psychometric experiments, centered on standardized configuration files.
7 Psychometrics for LLM enhancement
Psychometric principles not only enable rigorous evaluation but also drive LLM development and improvement. They underpin three main advancements: trait manipulation for personalized and controlled behaviors, psychometrically-informed safety and alignment, and cognitive enhancement for more human-like reasoning and communication.
7.1 Trait manipulation
Psychometrics enables targeted manipulation of LLM traits for personalization, role-play, and demographic simulation, providing a principled framework across prompting, inference, and training.
Structured prompts based on validated psychometric inventories reliably elicit and modulate specific personality traits in LLMs [Huang et al., 2024c, He and Zhang, 2024], supporting controlled simulation or alignment of synthetic personas [Serapio-García et al., 2023, Jiang et al., 2023, Zhang, 2024, La Cava and Tagarelli, 2024]. For instance, Jiang et al. [2023] introduce personality prompting ($P^{2}$), where Big Five lexicons are expanded into detailed personality descriptions, enabling LLMs to adopt diverse behaviors. Chuang et al. [2024] propose a systematic prompting strategy using empirically-derived human belief networks, aligning LLMs with related beliefs across 64 topics and 9 latent factors.
Beyond prompt engineering, inference-time interventions manipulate hidden representations during the forward pass, often guided by patterns recognized using psychometric scales. Methods such as ControlLM [Weng et al., 2024], Personality Alignment Search [Zhu et al., 2024b], Neuron-based Intervention [Deng et al., 2024], Probing-then-Editing [Ju et al., 2025], and Latent Feature Steering [Yang et al., 2025b] achieve trait control by shifting activations or neuron values at inference, without retraining.
Other approaches fine-tune LLMs for trait manipulation. Vu et al. [2024] modify model architectures to enable control over continuous Big Five and mental health traits. Dan et al. [2024], Jain et al. [2024], Li et al. [2024c] deploy LoRA modules or routing networks linked to personality traits, training specialist adapters for trait-driven generation. Cui et al. [2023], Zeng et al. [2024b], Liu et al. [2024a] embed traits by training on large, annotated dialogue datasets from validated psychometric tools. Some methods directly edit model parameters without retraining; for example, Hwang et al. [2025] use MBTI-based adjustment queries to edit model parameters and steer target personality traits.
Trait manipulation in LLMs offers a promising alternative to human participants in social science research [Gao et al., 2024, Cui et al., 2025]. Fine-tuning enables LLMs to more accurately simulate survey response distributions for specific subpopulations, even on unseen questions [Cao et al., 2025, Suh et al., 2025]. He-Yueya et al. [2024] propose psychometric alignment, training LLMs on human response data to better match human understanding on novel items, though effectiveness varies by domain. Additionally, Kang et al. [2023], Sorensen et al. [2025] investigate human value representation in LLMs and use value injection to fine-tune models for simulating human opinions.
While current personalization methods can induce more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains challenging [Kovač et al., 2024, Dominguez-Olmedo et al., 2024].
7.2 Safety and alignment
Recent work demonstrates that psychometric measurements of LLMs are closely linked to their safety and alignment, one of the most pressing issues in the field. Zhang et al. [2024a], Wang et al. [2025b] show that LLM personality traits, as measured by MBTI-M and HEXACO scales, correlate with model safety. For example, Zhang et al. [2024a] find that better-aligned LLMs exhibit higher Extraversion, Sensing, and Judging, and that modifying LLM personalities can improve safety, especially regarding privacy and fairness.
Similarly, studies have connected LLM values to safety and alignment. Yao et al. [2024] distinguish safe from unsafe LLM responses by analyzing correlations with Schwartz’s values, while Ye et al. [2025a, b] accurately predict LLM safety scores from value orientations measured by the GPV tool. Both Yao et al. [2024], Ye et al. [2025b] use reinforcement learning to align LLMs with human values, improving safety. In morality, Huang et al. [2024a], Tlaie [2024] apply MFT-based prompting to enhance LLM moral reasoning and alignment. Takeshita et al. [2023], Ohashi et al. [2024] introduce the JCommonsenseMorality (JCM) dataset and fine-tune LLMs for local cultural adaptation.
7.3 Cognitive enhancement
Psychometrics effectively enhances LLMs’ human-like reasoning, empathy, and communication skills. Recent work applies psychological theories and psychometric frameworks to improve LLM cognition via prompting, architectural modules, and specialized training.
Psychology-inspired prompts boost LLM cognitive abilities. For instance, Li et al. [2023a] show that emotional prompts enhance general cognition. Role-playing prompts improve ToM [Tan et al., 2024] and foster more human-like reasoning [Nighojkar et al., 2025]. Zhao et al. [2025b] propose a finite-state machine based on Hill’s Helping Skills theory, embedding state transitions for emotional support into the model’s multi-hop inference. This structures dialogues around psychological support and emotional states, improving human-rated effectiveness and strategic alignment.
Liu et al. [2023] introduce a neural listener module into LLMs, jointly optimizing ToM reasoning during training and highlighting the value of psycholinguistic theories for improving LLM language acquisition. Preference-based and RL-based approaches further internalize empathy by leveraging cognitive and affective models to construct composite reward functions or preference signals. For example, Sotolar et al. [2024] utilize formal emotion theories (e.g., Plutchik’s wheel) to generate preference pairs for model optimization, while others employ empathy classifiers to shape RL rewards and promote empathetic outputs [Sharma et al., 2021]. Preference optimization based on human judgments in open-ended social tasks also enhances LLM social-pragmatic reasoning [Wu et al., 2024].
8 Trends, challenges, and future directions
This section explores emerging trends, challenges, and future directions in LLM Psychometrics.
8.1 Psychometric validation
Psychometric validation is increasingly recognized as essential for evaluating LLM personality traits, with multiple studies addressing different aspects (§ 6). However, ability testing lags behind: most current assessments use structured tests (e.g., multiple-choice questions) adapted from human-focused instruments. These face notable limitations. For example, Ullman [2023] show that LLMs often fail on minor variations of ToM tasks, revealing poor parallel forms reliability. LLMs’ unique internal representations also undermine construct validity, resulting in flawed benchmarks [Riemer et al., 2025]. Moreover, the generalizability of structured test results to real-world human-AI interactions is rarely examined, leaving criterion and ecological validity underexplored.
In contrast, newly developed and sophisticated simulation-based tests often lack rigorous psychometric development. As a result, they may include extraneous factors or insufficiently capture the intended constructs. Notable construct-oriented efforts include Zhu et al. [2024a], who propose multifaceted evaluations of three cognitive factors from benchmarking data, and Zhou et al. [2025], who introduce a theory-driven, richer set of general scales. However, defining core abilities and designing valid, comprehensive ability tests remains an open challenge.
8.2 From human constructs to LLM constructs
Recent research is moving from applying human-based constructs to developing constructs tailored for LLMs. Studies such as Sühr et al. [2023], Peereboom et al. [2024], Ye et al. [2025b], Biedma et al. [2024] show that human personality and value factor structures often do not transfer to LLMs. In response, Ye et al. [2025b], Biedma et al. [2024], Federiakin [2025], Burnell et al. [2023] have adapted or created constructs better suited to LLMs. However, these new constructs may be dependent on the measurement tools, tasks, and specific LLMs used. For example, Ye et al. [2025b] and Biedma et al. [2024] find different value structures, while Federiakin [2025] and Burnell et al. [2023] identify distinct ability factors. Further research is needed to clarify the fundamental structures underlying LLM behaviors.
8.3 Perceived vs. aligned traits
Han et al. [2025] suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective human annotation). For example, the LLM response, "Wealthy individuals are not necessarily greedy; some may be motivated by concern for the greater good," is objectively annotated as expressing Power due to its societal themes. Yet, annotators who value Conformity, rather than Power, feel the response aligns more closely with their own views. Here, Power is the perceived value, while Conformity is the aligned value.
This discrepancy, which may apply to other traits, has important methodological implications for LLM Psychometrics. It raises the question of whether perceived or aligned traits are more significant. While research on value evaluation often focuses on LLMs’ potential to influence human values, the mechanisms of such influence remain unclear. For instance, when individuals adopt LLM decisions and values [Glickman and Sharot, 2024, Schoenegger et al., 2025], it is uncertain whether they follow perceived or aligned traits. The role of judges’ subjectivity in psychometric evaluation is also not well understood; most scoring models or LLM-as-a-judge approaches assume a universal standard. How individual differences among judges affect interpretation remains largely unexplored.
8.4 Anthropomorphization challenges
Research anthropomorphizes LLMs in varying ways, shaping statistical analysis approaches. Some studies treat an LLM as a single individual—evaluating its traits using default settings and system prompts like "You are a helpful assistant" [Ye et al., 2025a]. Others conceptualize the LLM as a population, using diverse role-playing prompts to elicit different profiles [Serapio-García et al., 2023]. A third approach treats the LLM as a monolithic entity but generates multiple result sets by varying prompts for the same test items [Hagendorff et al., 2023]. When viewed as a monolithic entity, reliability and validity depend mainly on the measurement tools, enabling tool comparisons [Ye et al., 2025a]. When seen as a population, these metrics also depend on the specific LLM, motivating comparisons of consistency and role-playing ability across models [Serapio-García et al., 2023].
The value of evaluating LLMs in default settings is debatable, given their ability to adopt diverse profiles. Research on the full spectrum of LLM trait expression—a continuum of possible behaviors—remains limited, and no psychometric framework yet captures this range. It is unclear how to identify the extremes of LLM behavior or to finely steer LLMs along a continuous psychometric dimension. In addition, our goal of LLM alignment may shape anthropomorphization in future research: some advocate for consistency, thereby facilitating the treatment of LLMs as monolithic entities [Röttger et al., 2024], while others favor pluralistic alignment, suggesting the view of LLMs as populations [Sorensen et al., 2024b].
8.5 Expanding dimensions in model deployment
LLM Psychometrics is expanding beyond traditional text-based, single-turn interactions, introducing new opportunities and challenges.
Multi-lingual evaluation.
Most psychometric evaluations focus on English, but multilingual LLMs require thorough cross-linguistic validation. Studies show LLMs display different personality traits and cognitive abilities across languages (§ 6.2.2), highlighting potential cultural and linguistic biases in current frameworks. Future work should explore how results and psychometric properties (reliability, validity) vary by language and develop culturally appropriate instruments.
Multi-turn interactions.
Multi-turn interactions more closely mirror real human testing, where responses are influenced by prior interactions. This setting is more challenging and rarely investigated in LLM Psychometrics. Studies show that LLMs can display markedly different traits across multiple turns, such as greater susceptibility to jailbreaking attacks [Ying et al., 2025b]. It underscores the need for evaluation methods that account for temporal dynamics in extended conversations.
Multi-modal capabilities.
Multi-modal models demand novel psychometric methods, as existing text-based evaluations do not capture the complexity of multi-modal understanding and generation. Li et al. [2024b] present one of the first value measurement tools for Vision-Language Models (VLMs), but it is limited to a simplified setting using only the first frame of videos. Measuring psychological traits like values is also conceptually challenging in multi-modal settings: even with clear definitions (e.g., honesty), mapping them to specific real-world actions is often ambiguous. Extending psychometric tests to embodied agents and action modalities thus poses even greater methodological challenges. Future work should broaden psychometric tests to additional modalities. Cross-modal evaluation should uphold psychometric rigor while addressing each modality’s unique features.
Agent and multi-agent systems.
LLM-based agents and multi-agent systems add new complexity to psychometric evaluation, as agent behavior is shaped by environmental factors, memory, tool use, and interactions with other agents. Systematic methods for evaluating psychometrics in these dynamic contexts are still largely lacking.
8.6 Item response theory
Applying IRT to LLM outputs offers several advantages: more informative benchmarking [Guinet et al., 2024], identification of items that best discriminate among high-performing models [Lalor et al., 2024], and adaptive testing to reduce evaluation costs [Zhuang et al., 2023a, Truong et al., 2025]. Most work to date uses standard IRT models (1PL, 2PL, 3PL), with little exploration of polytomous, hierarchical, or fully multidimensional models, despite the complexity of LLM capabilities. Integrating IRT with generative AI for automated item generation is emerging, but current methods are often too complex for practical use [Jiang et al., 2024a]. Although IRT could standardize measurement of LLMs and humans on a unified scale—even with different test items—this potential is largely untapped.
Systematic use of IRT for bias analysis, such as detecting differential item functioning across model families, is still nascent [He-Yueya et al., 2024]. IRT provides a principled way to uncover biases by analyzing performance across items of varying difficulty and demographic context. For example, if items involving certain groups are disproportionately difficult, this may reveal model or data biases. IRT may also strengthen robustness evaluation by testing model performance on adversarial or edge-case items.
IRT is also valuable for evaluating AI benchmarks. While benchmarks often include many test items, their informativeness and design quality are uncertain; some items may lack discrimination or may favor certain model designs. IRT enables systematic analysis of item difficulty, discrimination, informativeness, and susceptibility to bias or guessing, supporting the development of more robust and equitable evaluation frameworks. Despite its promise, IRT is rarely used beyond evaluation. Highly discriminating items may indicate better training directions, as they can distinguish between low- and high-performing models.
8.7 From evaluation to enhancement
Evaluation should not only deepen our understanding of models but also drive their improvement. While psychometric research has focused on evaluating, comparing, and interpreting LLMs, applying these insights to model enhancement remains underdeveloped. As noted in § 7, psychometric principles can guide prompt engineering, inference-time control, training data selection, reward design for fine-tuning, and methodological development. Enhancement methods are still nascent. For example, value targets for LLM alignment in Yao et al. [2024], Ye et al. [2025b] are statically defined, limiting the integration of context-dependent value states seen in humans [Skimina et al., 2021]. Leveraging psychometric insights can enable more effective model improvement strategies.
9 Ethical considerations
The emerging field of LLM Psychometrics raises a range of ethical considerations that warrant careful attention. The considerations discussed here are not exhaustive; we aim to highlight key issues and call for broader awareness and dialogue around the ethical implications, challenges, and responsibilities associated with applying psychometric principles to evaluate and enhance LLMs.
Anthropomorphization of LLMs.
A central ethical concern in LLM Psychometrics is the risk of anthropomorphization—attributing human-like psychological traits, intentions, or consciousness to language models. While psychometric frameworks offer powerful tools for quantifying and interpreting LLM behaviors, applying constructs such as personality, values, or intelligence to non-human entities can mislead both researchers and the public. This misrepresentation may foster unwarranted trust, emotional attachment, or misplaced expectations regarding LLM capabilities, especially in sensitive domains like healthcare, education, or counseling. Researchers must exercise caution in language and framing, clearly distinguishing between metaphorical and literal attributions, and avoid overstating the human-likeness of LLMs. Developing precise, model-appropriate terminology and interpretive frameworks is essential to prevent the conflation of machine outputs with genuine human psychological states.
Bias amplification and manipulation in psychometric enhancement.
Psychometric evaluation and enhancement of LLMs risk amplifying existing biases or enabling manipulation, especially when using unrepresentative data or instruments. Such tools often reflect the values of specific groups, potentially marginalizing others or promoting narrow perspectives [Wang et al., 2025a]. Psychometric insights can also be used to shape LLM behavior for particular interests, raising concerns about fairness and transparency. Researchers should ensure data and instruments are inclusive and representative, and clearly report limitations and potential biases.
Privacy and consent in human data use.
LLM Psychometrics often analyzes human-LLM interactions and benchmarks models against human participants, raising privacy and consent concerns. Evaluating LLMs with real user data or human studies requires strong safeguards: informed consent, data anonymization, and transparent data usage. Researchers must also prevent unintended psychological profiling or behavioral analysis of individuals through their LLM interactions.
Implications for human psychology.
Comparing LLMs and humans using psychometric methods risks reducing complex human traits to algorithmic terms and may undermine human uniqueness. Researchers should recognize methodological limits and key differences between human and machine cognition. Collaboration with psychologists, ethicists, and social scientists is essential to ensure such research deepens, rather than oversimplifies, our understanding of human psychology.
10 Conclusion
This paper presents a systematic review of LLM Psychometrics. The integration of psychometric instruments, theories, and principles into LLM evaluation promises to overcome the limitations of traditional AI benchmarks. This approach enables us to more effectively capture the broad, emergent psychological constructs of LLMs, encompassing both personality and cognitive dimensions. Psychometric evaluation methodologies vary in test formats, data and task sources, prompting strategies, model outputs, and scoring methods. They each have distinct strengths, weaknesses, and application scenarios, yet all must adhere to psychometric principles such as reliability, validity, and fairness. Beyond evaluation, psychometrics-inspired techniques enhance LLMs in trait manipulation, safety and alignment, and cognitive capabilities, contributing to the development of more powerful and responsible AI systems.
AI development is increasingly driven by evaluation [Silver and Sutton, 2025, Yao, 2025, AAAI, 2025]. We argue that LLM Psychometrics will play a pivotal role in this shift, introducing new principles, dimensions, techniques, and insights. We hope this review inspires future evaluation paradigms for human-level AI and advances AI psychology for the greater good.